### v1

# Totally elliptic surface group representations

Abstract.
We characterize all totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ by showing that they are either representations into a compact subgroup or Deroin–Tholozan representations.
1. Introduction
1.1. Main result
Surface group representations are homomorphisms from the fundamental group of a surface $\Sigma$ into a Lie group $G$. Their deformation spaces are called character varieties. A character variety is not necessarily a connected space. Dynamical behaviour and qualitative properties of representations may vary considerably between components. At one end of the spectrum are the higher Teichmüller spaces which are a generalization of the classical and extensively studied Teichmüller space of $\Sigma$. Higher Teichmüller spaces, as defined in [Wie18], consist of discrete and faithful representations and tend to behave like moduli spaces of geometric structures on $\Sigma$, although a precise correspondence is still unavailable. This article aims at drawing the reader’s attention towards the other end of the spectrum where connected components are mostly made of non-discrete representations.
We’ll focus on a particular prototype of representations that are antipodal to discrete and faithful ones: the so-called totally elliptic representations.
Definition 1.1.
A surface group representation $\pi_{1}\Sigma\to G$ is totally elliptic if every simple closed curve on $\Sigma$ is mapped to an elliptic element of $G$. By an elliptic element of $G$, we mean an element of a compact subgroup of $G$ [^1] Those elements are sometimes called compact in the literature.
The surface $\Sigma$ will always be taken to be oriented and connected, and can be punctured. Recall that a simple closed curve on $\Sigma$ is the free isotopy class of a closed loop on $\Sigma$ that doesn’t self-intersect and isn’t homotopically trivial. A simple closed curve can be homotopic to a puncture, in which case we call it peripheral. By definition, the images of every peripheral curve by a totally elliptic representation are elliptic. We want to avoid the situation where peripheral curves are mapped to the identity and we say that a representation is reduced when this isn’t the case (Definition 2.7).
The first instances of totally elliptic representations, though not the most interesting, are those into a compact subgroup of $G$. Interestingly, however, by requiring only simple closed curves to be mapped to elliptic elements (as opposed to all curves), one can construct Zariski dense examples of totally elliptic representations into non-compact Lie groups, such as every Hermitian Lie group (Section 1.3), whose deformations remain totally elliptic.
The initial case to investigate is when $G$ is $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. In that case, the first totally elliptic representations with Zariski dense image were discovered by Benedetto–Goldman [BG99], and later generalized by Deroin–Tholozan [DT19]. Those representations were originally named supra-maximal in [DT19], but we’ll refer to them as DT representations (Definition 2.10), to avoid any confusion with maximal representations from higher Teichmüller theory. DT representations only exist when $\Sigma$ is a genus-0 surface with at least three punctures. As Deroin–Tholozan observed, their deformation spaces are compact components inside the corresponding relative character varieties and made exclusively of totally elliptic representations (Theorem 2.9). The question was raised whether they’re the only totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, besides representations into compact subgroups. We bring a positive answer to that question.
Theorem A.
Let $\Sigma$ be a connected and oriented surface with genus $g\geq 0$ and $n\geq 0$ punctures. Reduced totally elliptic representations $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ can only be of two kinds:
  • The image of $\rho$ is contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. In that case, we say that $\rho$ is an orthogonal representation.
  • If $\rho$ is not orthogonal, then $g=0$, $n\geq 3$, and $\rho$ is a DT representation.
We prove Theorem A in Section 2.
1.2. Motivations
Characterizing a certain type of surface group representations (such as totally elliptic ones) by the images of simple closed curves may seem baroque at first. Similar problems, however, have sparked substantial interest in recent decades. Bowditch famously raised the question of whether discrete and faithful representations of a closed surface group $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—also known as Fuchsian representations—can be characterized by the property of mapping every simple closed curve on $\Sigma$ to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (an element whose trace is larger than $2$ in absolute value) [Bow98, Question C]. If true, this would be a finer characterization of Fuchsian representations which are known to be the only representations that map every non-trivial element of $\pi_{1}\Sigma$ to hyperbolic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Bowditch’s question has been answered positively for surfaces of genus $2$ by Marché–Wolff [MW16, MW19] and remains open in higher genera.
Of course, not every representation is Fuchsian. Goldman proved that the character variety of representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma$ is a closed surface of genus $g\geq 2$, has $34g-$ 3 connected components [Gol88]. Two of the components are made of conjugacy classes of Fuchsian representations, forming two copies of the Teichmüller space of $\Sigma$; the other $54g-$ 5 components are known as intermediate components and they don’t contain any Fuchsian representation. Bowditch’s question is equivalent to asking whether every representation whose conjugacy class lies in an intermediate component maps at least one simple closed curve on $\Sigma$ to a non-hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Marché–Wolff also proved that a positive answer to Bowditch’s question would imply that the mapping class group action on every intermediate component is ergodic [MW16, Theorem 1.6]. The latter statement is known as Goldman’s ergodicity conjecture [Gol06, Conjecture 3.1] and is open, like Bowditch’s question, for surfaces of genus at least $3$.
Goldman also conjectured that almost every non-Fuchsian closed surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ should be the holonomy of a branched hyperbolic structure on $\Sigma$ which is the data of hyperbolic metric of $\Sigma$ with several conical singularities of angle $2\pi k$ with $k\geq 2$. To be precise, according to Goldman, every dense representation whose conjugacy class lies in an intermediate component should be the holonomy of a branched hyperbolic structure [Gol06, Conjecture 3.9]. Mathews proved some partial implication from a positive answer to Bowditch’s question to Goldman’s geometrization conjecture for intermediate components of co-Euler class $1$ [Mat12]. Observe as well that Goldman’s ergodicity conjecture implies a “full measure version” of the geometrization conjecture, as well as a positive answer to Bowditch’s question on a full measure subset of intermediate components.
Bowditch’s question Geometrization conjecture Ergodicity conjecture [MW16]
Goldman’s two conjectures and Bowditch’s question, and the interplay among them, constitute a whole program to study intermediate components which we refer to as the Bowditch–Goldman program.
Totally elliptic representations, and more particularly DT representations, naturally arise as a prototype of surface group representations that fit perfectly in the Bowditch–Goldman program. By definition, they map every simple closed curve to an elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, and therefore satisfy the analogue of Bowditch’s condition on intermediate representations in a strong sense. The ergodicity of the pure mapping class group action on the character variety of DT representations has been established in [Mar22]. As explained in [DT19, Section 4] and further detailed in [FM23], DT representations are also holonomies of hyperbolic cone structures with branch point singularities, thereby satisfying a precise version of the geometrization conjecture.
1.3. Perspectives away from $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$
This note is also an opportunity to survey some recent developments about totally elliptic representations into Lie groups $G$ different than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Several families of compact connected components of totally elliptic representations were identified when $G$ is a Hermitian Lie group. They’re all genus-0 surface group representations. The first examples were found by Tholozan–Toulisse when $G$ is the Lie group $\mathrm{SU}(p,q)$, as well as when $G$ is $\mathrm{Sp}(2n,\operatorname{\mathbb{R}})$ or $\mathrm{SO}^{\ast}(2n)$ [TT21]. Analogous compact components have been identified by Feng–Zhang when $G$ is the identity component of $\mathrm{SO}(2,n)$ [FZ23]. Their methods studies the topology of some moduli spaces of parabolic Higgs bundles which relate to relative character varieties via the non-abelian Hodge correspondence. In particular, they proved that for every representation $\rho\colon\pi_{1}\Sigma\to G$ whose conjugacy class belongs to one of the compact components they identified, and for every punctured Riemann sphere structure on $\Sigma$, there exists a $\rho$-equivariant holomorphic map from the universal cover $\widetilde{\Sigma}$ to the symmetric space $G/K$. The same property for DT representations was proven by Mondello [Mon16]. Surface group representations satisfying this condition are sometimes called universal variations of Hodge structures.
Unlike DT representations which are all Zariski dense, it’s not clear yet whether every compact component of totally elliptic representations into Hermitian Lie groups other than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ contains a Zariski dense representation (almost all of them do though, by the preponderance of Zariski dense representations in any character variety). The topology of those compact components is also yet to be determined, whereas DT components are known to be homeomorphic (even symplectomorphic) to complex projective spaces [DT19, Theorem 4].
There are however more explicit examples of compact components of totally elliptic representations discovered by Marc-Zwecker [MZ24]. Marc-Zwecker’s examples are made of representations of a pair of pants into $\mathrm{SU}(2,1)$ where the peripheral curves are mapped to products of reflections through three complex geodesic lines in the complex hyperbolic space of real dimension four. The three complex geodesic lines are the mirrors and they’re arranged as a triangle. The resulting components are homeomorphic to 2-dimensional spheres. The geometric nature of the representations in Marc-Zwecker’s components may lead to other compact components for spheres with a higher number of punctures obtained by “chaining” the triangles of mirrors in the spirit of [Mar24].
A natural question to ask is whether the class of Lie groups for which there exist substantial connected components entirely made of totally elliptic genus-0 surface group representations is larger than the class of Hermitian Lie groups.
Question 1.
Are there semisimple real Lie groups other than Hermitian Lie groups for which there exist compact connected components (of maximal expected dimension) inside relative character varieties of genus-0 surface group representations made of totally elliptic representations and containing Zariski dense representations?
There are reasons to believe that the key property is the presence of a nonempty open subset of elliptic elements in the target Lie group. This property is shared by all Hermitian Lie groups, but not only; for instance, Sugiura proved that $\mathrm{SO}(2p,q)$ has an open subset of elliptic elements [Sug59, Theorem 8], even though it’s not Hermitian if $p\geq 2$ and $q\geq 3$. In contrast, the subset of elliptic elements in $\mathrm{SL}_{n}\operatorname{\mathbb{R}}$ has empty interior for every $n\geq 3$. Furthermore, Kabenyuk proved that admitting a nonempty open subset of elliptic elements is equivalent to having a compact Cartan subgroup [Kab90].
One should nevertheless be aware that “large” Lie groups, unlike $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$, may admit subgroups $\Gamma$ containing only elliptic elements but such that $\Gamma$ is not contained into a compact subgroup. Bass gives an example of such a $\Gamma$ which is isomorphic to a rank-2 free subgroup of $\mathrm{SL}_{3}\operatorname{\mathbb{C}}$ [Bas80, Counterexample 1.10]. Bass’ example is not Zariski dense and, in fact, it couldn’t be as explained for instance in [Pra94]. However, when $\operatorname{\mathbb{C}}$ is replaced by a non-Archimedean valued field $\mathbb{F}$, then the situation is similar to the case of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ as Parreau proved that a finitely generated subgroup of $\mathrm{SL}_{n}\mathbb{F}$ containing only elliptic elements is necessarily bounded [Par03, Théorème 1]. This doesn’t rule out the existence of interesting totally elliptic surface group representations into $\mathrm{SL}_{n}\mathbb{F}$.
1.4. Acknowledgments
Many thanks to Samuel Bronstein and Nicolas Tholozan for enriching conversations on totally elliptic representations. I’m grateful to Bertrand Deroin for sparking my curiosity about the conjectures related to the Bowditch–Goldman program, and to Maxime Wolff for teaching me the elegant geometric argument used to prove Lemma 2.1. I also extend my gratitude to my friends Jacques Audibert and Xenia Flamm for their insightful comments on an earlier draft of this paper and for pointing out to me Parreau’s paper.
2. Proof of Theorem A
2.1. Overview
The proof of Theorem A unfolds over the next subsections. We start by showing that when the genus of $\Sigma$ is at least $1$, then the only totally elliptic representations are orthogonal representations (Section 2.2). The arguments mostly rely on the observation that a commutator of elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is itself elliptic only when it’s trivial (Lemma 2.1). We then briefly recall what DT representations are (Section 2.3) and prove that a totally elliptic genus-0 surface group representation that isn’t orthogonal must be a DT representation (Sections 2.3.12.3.3).
2.2. Positive genus
We start by proving that a totally elliptic representation $\rho\colon\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of a surface $\Sigma$ of genus $g\geq 1$ is necessarily an orthogonal representation (Propositions 2.5 and 2.8). Most of the reasoning will rely on the following standard results about $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we include for the sake of completeness.
According to Definition 1.1, the elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ fall into two categories: the regular elliptic elements and the identity. The regular elliptic elements are those whose trace is smaller than $2$ in absolute value. Geometrically, they act on the hyperbolic plane $\operatorname{\mathbb{H}}$ by rotations around their unique fixed point. The subset of regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is homeomorphic to the open ball $(0,2\pi)\times\operatorname{\mathbb{H}}$, where the first factor records the rotation angle of a regular elliptic and the second its unique fixed point (the centre of rotation).
Lemma 2.1.
A regular elliptic element $A$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is contained in a unique maximal compact subgroup which consists of all the elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ which fix the unique fixed point of $A$. Furthermore, if $X$ is any element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, then the commutator
$[A,X]=AXA^{-1}X^{-1}$
is either the identity or a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. The commutator is trivial if and only if $X$ is elliptic and fixes the unique fixed point of $A$.
Proof.
The unique maximal compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ that contains $A$ is its centralizer, which is conjugate to $\operatorname{PSO}(2)$. Since $A$ is regular, any element that commutes with $A$ is either the identity or a regular elliptic with the same fixed point as $A$.
Let’s now consider the commutator $[A,X]$. As we just said, if $[A,X]=1$, then $X$ is elliptic and fixes the unique fixed point of $A$. If $[A,X]\neq 1$ and $o$ denotes the unique fixed point of $A$, then $XA^{-1}X^{-1}$ is regular elliptic with fixed point $X.o\neq o$. Note that if $\alpha$ is the rotation angle of $A$, then the rotation angle of $XA^{-1}X^{-1}$ is $2\pi-\alpha$. Let $\ell$ denote the unique geodesic line in $\operatorname{\mathbb{H}}$ through $o$ and $X.o$. Similarly, let $\ell_{1}$ and $\ell_{2}$ denote the geodesic lines obtained by rotating $\ell$ anti-clockwise by an angle $\alpha/2$ around $o$, respectively $X.o$.
[…]
The orientation-reversing reflections of $\operatorname{\mathbb{H}}$ through the lines $\ell$, $\ell_{1}$, and $\ell_{2}$ are denoted by $\sigma$, $\sigma_{1}$, and $\sigma_{2}$, respectively. With this notation, we can write $A=\sigma_{1}\sigma$ and $XA^{-1}X^{-1}=\sigma\sigma_{2}$, so that $[A,X]=\sigma_{1}\sigma_{2}$. We conclude that $[A,X]$ preserves the common orthogonal geodesic line to $\ell_{1}$ and $\ell_{2}$, and is therefore hyperbolic. ∎
Lemma 2.1 plays a central role in our discussion because it tells us that the commutator of two elliptic elements is elliptic if and only if it is trivial. It turns out that the commutator of two simple closed curves (seen as elements of $\pi_{1}\Sigma$) sometimes represents a simple closed curve too. We must, however, be careful, because a simple closed curve on $\Sigma$ is represented by many elements in $\pi_{1}\Sigma$, all conjugate to each other.
Lemma 2.2.
If two simple closed curves on $\Sigma$ intersect once (their geometric intersection number is 1) and we represent them in the simplest way by two fundamental group elements $a$ and $b$ with basepoint being the intersection point, then $[a,b]$ also represents a simple closed curve on $\Sigma$.
Proof.
Pick two loops on $\Sigma$ that only intersect at the basepoint and whose homotopy classes are $a$ and $b$. Enlarge each loop by replacing it with a thin ribbon. The boundary of the union of the two ribbons is an embedded circle in $\Sigma$ which is in the free homotopy class determined by the fundamental group element $[a,b]$. ∎
As we’re about to see, Lemmas 2.1 and 2.2 together are the main obstruction for constructing totally elliptic surface group representations. Let’s start by introducing some notation. We pick a geometric generating family for $\pi_{1}\Sigma$ which consists in $2g+n$ generators $a_{1},\ldots,a_{g}$, $b_{1},\ldots,b_{g}$, and $c_{1},\ldots,c_{n}$ satisfying the unique relation
$\prod_{i=1}^{g}[a_{i},b_{i}]=\prod_{i=1}^{n}c_{i}.$
A possible choice of geometric generators are the homotopy classes of the loops illustrated on Figure 1.
Figure 1. A system of geometric generators for a surface $\Sigma$ of genus $g$ with $n$ punctures.
Remark 2.3.
We adopt the convention of concatenating loops by starting with the rightmost loop. More precisely, if $\gamma_{1}$ and $\gamma_{2}$ are two loops on $\Sigma$ based at the same point, then their concatenation $\gamma_{1}\gamma_{2}$ is the loop obtained by going along $\gamma_{2}$ first, and then along $\gamma_{1}$. This rule applies to fundamental group multiplications too.
Lemma 2.4.
If the genus $g$ of $\Sigma$ is at least 2 and $i\neq j$ are two indices in $\{1,\ldots,g\}$, then the following fundamental group elements represent simple closed curves on $\Sigma$:
$[a_{i},b_{i}],\quad[a_{j},b_{j}],\quad[b_{i}^{-1}b_{j}^{-1},a_{i}],\quad[a_{i}^{-1}a_{j}^{-1},b_{j}],$
$[b_{j}^{-1}b_{i}^{-1},a_{j}],\quad[a_{j}^{-1}a_{i}^{-1},b_{j}],\quad[b_{i}a_{i}^{-1}b_{j},a_{j}],\quad[a_{i}b_{i}^{-1}a_{j},b_{j}].$
Proof.
Each pair of fundamental group elements appearing in a given commutator represents two simple closed curves on $\Sigma$ with geometric intersection number $1$. So, by Lemma 2.2, we can expect their commutators to represent simple closed curves too. We must, however, be careful when we represent a simple closed curve by a fundamental group element because of basepoint issues. The choice we made ensures that each commutator indeed represents a simple closed curve on $\Sigma$. This can be seen by drawing the corresponding curves on Figure 1. ∎
Proposition 2.5.
Assume that $\Sigma$ is closed and has genus $g\geq 1$. If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a totally elliptic representation, then $\rho$ is an orthogonal representation.
Proof.
We’ll write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$. If $g=1$, then the image of $\rho$ is generated by $A_{1}$ and $B_{1}$ which satisfy $[A_{1},B_{1}]=1$. So, the image of $\rho$ is generated by two commuting elliptic elements, hence contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
Let’s now consider the case $g\geq 2$. If $\rho$ is the trivial representation, then it’s orthogonal. If not, then there exists an index $i\in\{1,\ldots,g\}$ such that either $A_{i}\neq 1$ or $B_{i}\neq 1$. We’ll look at the case where $A_{i}\neq 1$ (the case where $B_{i}\neq 1$ is identical). Let $j\neq i$ be another index. Because $\rho$ is totally elliptic, it maps all the simple closed curves from Lemma 2.4 to the identity because of Lemma 2.1. In particular, we have $[A_{i},B_{i}]=1$ and $[A_{i},B_{i}^{-1}B_{j}^{-1}]=1$, implying that $[A_{i},B_{j}]=1$. Since $[A_{i},B_{i}]=1$, up to conjugating $\rho$, we may assume that $A_{i}$ and $B_{i}$ belong to $\operatorname{PSO}(2)$. From $[A_{i},B_{j}]=1$ and our assumption that $A_{i}\neq 1$, we conclude that $B_{j}\in\operatorname{PSO}(2)$ by Lemma 2.1.
We also infer from Lemmas 2.1 and 2.4 that $[A_{j},B_{j}]=[A_{j},B_{j}^{-1}B_{i}^{-1}]=[A_{j},B_{i}A_{i}^{-1}B_{j}]=1$. If $B_{j}\neq 1$, then we immediately conclude that $A_{j}\in\operatorname{PSO}(2)$ from $[A_{j},B_{j}]=1$ by Lemma 2.1. Since $[A_{j},B_{j}]=1$ and $[A_{j},B_{j}^{-1}B_{i}^{-1}]=1$, we deduce that $[A_{j},B_{i}]=1$ which also implies $A_{j}\in\operatorname{PSO}(2)$ as soon as $B_{i}\neq 1$. If both $B_{i}=1$ and $B_{j}=1$, then we alternatively observe that $[A_{i}^{-1},A_{j}]=[B_{i}A_{i}^{-1}B_{j},A_{j}]=1$ which gives $A_{j}\in\operatorname{PSO}(2)$ as well, because $A_{i}\neq 1$.
In conclusion, we proved that $A_{i},B_{i},A_{j},B_{j}\in\operatorname{PSO}(2)$. Since the indices $i$ and $j$ were picked arbitrarily, we conclude that the image of $\rho$ is contained in $\operatorname{PSO}(2)$, proving that $\rho$ is orthogonal. ∎
We now consider the case where $\Sigma$ has punctures. We first identify more simple closed curves.
Lemma 2.6.
If $\Sigma$ is a surface of genus $g\geq 1$ and with $n\geq 1$ punctures, then for any index $i\in\{1,\ldots,g\}$ and $j\in\{1,\ldots,n\}$ the following fundamental group elements represent simple closed curves on $\Sigma$:
(2.1) $a_{i}^{-1}[a_{i},b_{i}]c_{j}^{-1}a_{i}c_{j}\quad\text{and}\quad b_{i}[a_{i},b_{i}]c_{j}^{-1}b_{i}^{-1}c_{j}.$
If furthermore $n\geq 2$ and $k\in\{1,\ldots,n\}$ is another index such that $j>k$, then
(2.2) $b_{i}c_{k}b_{i}^{-1}c_{j}b_{i}a_{i}^{-1}b_{i}^{-1}c_{k}^{-1}b_{i}^{-1}c_{j}^{-1}$
represents a simple closed curve too.
Proof.
The first fundamental group element in (2.1) can be rewritten as $b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}a_{i}c_{j}$, which is the product of a conjugate of $a_{i}^{-1}$ and a conjugate of $a_{i}$. A similar observation holds for the second fundamental group element in (2.1) which is equal to $b_{i}a_{i}b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}b_{i}^{-1}c_{j}$ and is thus a product of conjugates of $b_{i}$ and $b_{i}^{-1}$. By drawing a rapid picture, we convince ourselves that both represent simple closed curves on $\Sigma$. We prove in the same way that the fundamental group element (2.2) represents a simple closed curve too. ∎
Definition 2.7.
A surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is said to be reduced if it doesn’t map any peripheral curve on $\Sigma$ to the identity. In terms of the geometric generators from Figure 1, a representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is reduced if it doesn’t map any generator $c_{i}$ to the identity.
Note that every non-reduced totally elliptic surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ corresponds to a reduced totally elliptic representation $\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma^{\prime}$ is obtained from $\Sigma$ by removing the punctures whose peripheral curves are mapped to the identity.
Proposition 2.8.
Assume that $\Sigma$ is a surface of genus $g\geq 1$ and with $n\geq 1$ punctures. If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a reduced totally elliptic representation, then $\rho$ is an orthogonal representation.
Proof.
As before, we’ll write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$, as well as $C_{j}=\rho(c_{j})$. Proposition 2.5 implies that all the $A_{i}$ and $B_{i}$ are contained in the same orthogonal subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we can assume to be $\operatorname{PSO}(2)$ up to conjugating $\rho$. In particular, $[A_{i},B_{i}]=1$ for every index $i$.
Let’s first assume that there exists an index $i\in\{1,\ldots,g\}$ such that $A_{i}\neq 1$ or $B_{i}\neq 1$. We’ll consider the case where $A_{i}\neq 1$ (similar arguments apply when $B_{i}\neq 1$). Using the simple closed curves (2.1) from Lemma 2.6 and since $\rho$ is totally elliptic, we conclude that $A_{i}^{-1}[A_{i},B_{i}]C_{j}^{-1}A_{i}C_{j}$ is elliptic for every index $j$. Because $[A_{i},B_{i}]=1$, we even have that $[A_{i}^{-1},C_{j}^{-1}]$ is elliptic. As both $A_{i}$ and $C_{j}$ are elliptic by total ellipticity of $\rho$, Lemma 2.1 implies that $[A_{i},C_{j}]=1$ and thus $C_{j}\in\operatorname{PSO}(2)$ because $A_{i}\neq 1$. This holds for every index $j$ and therefore the image of $\rho$ is contained in $\operatorname{PSO}(2)$.
Now, let’s consider the case where $A_{i}=B_{i}=1$ for every $i\in\{1,\ldots,g\}$. Since we’re assuming $\rho$ to be reduced, we must have $n\geq 2$. Using the simple closed curve (2.2) from Lemma 2.6, we deduce that $B_{i}C_{k}B_{i}^{-1}C_{j}B_{i}A_{i}^{-1}B_{i}^{-1}C_{k}^{-1}B_{i}^{-1}C_{j}^{-1}$ is elliptic for every index $i$ and $j>k$. Since $A_{i}=B_{i}=1$, we even have $[C_{k},C_{j}]=1$ for every $j>k$ in $\{1,\ldots,n\}$. This implies that $\rho$ is orthogonal for the same reasons as before. ∎
Propositions 2.5 and 2.8 say that the only examples of reduced totally elliptic representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ when $\Sigma$ has positive genus and an arbitrary number of punctures are orthogonal representations.
2.3. Genus zero
In order to finish the proof of Theorem A, we shall prove that a reduced totally elliptic genus-0 surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is either orthogonal or a DT representation. Clearly, if $\Sigma$ has at most two punctures, then $\rho$ is orthogonal. Let’s now assume that $\Sigma$ is a sphere with a set $\mathcal{P}$ of $n\geq 3$ punctures. Recall that the simple closed curves on $\Sigma$ that are homotopic to a puncture are called peripheral. By definition, a reduced totally elliptic representation maps every peripheral curve on $\Sigma$ to a regular elliptic element. In other words, the conjugacy class $[\rho]$ of a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ belongs to an $\alpha$-relative character variety
$\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$
which consists of all conjugacy classes of representations mapping peripheral curves around every puncture $p\in\mathcal{P}$ to regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of fixed rotation angle $\alpha_{p}\in(0,2\pi)$. The numbers $\alpha_{p}$ are the components of the angle vector $\alpha\in(0,2\pi)^{\mathcal{P}}$. We’ll write
$|\alpha|=\sum_{p\in\mathcal{P}}\alpha_{p}.$
Unlike in the positive genus case, there are Zariski dense totally elliptic representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ when $\Sigma$ is an $n$-punctured sphere. The first examples were discovered via computer simulations by Benedetto–Goldman in the case $n=4$. Their construction was later generalized by Deroin–Tholozan for spheres with an arbitrary number of punctures.
Theorem 2.9 ([BG99, DT19]).
If $|\alpha|<2\pi$ or $|\alpha|>2\pi(n-1)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains a compact connected component isomorphic to $\mathbb{CP}^{n-3}$ and entirely made of conjugacy classes of Zariski dense totally elliptic representations that don’t contain any simple closed curve in their kernel.
Definition 2.10.
A representation whose conjugacy class belongs to one of the compact components of Theorem 2.9 will be called a DT representation. The components themselves will be referred to as DT components.
According to Theorem 2.9, DT representations are more than just totally elliptic in the sense of Definition 1.1; they are “regularly totally elliptic” by which we mean that every simple closed curve is mapped to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
DT representations can be characterized by their Toledo number in the sense of Burger–Iozzi–Wienhard [BIW10]. Namely, according to Deroin–Tholozan’s original definition, a conjugacy class of representations in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ belongs to a DT component if and only if its Toledo number is
(2.3) $\left\{\begin{array}[]{ll}1-\frac{|\alpha|}{2\pi},&\text{when $|\alpha|<2\pi$},\\ (n-1)-\frac{|\alpha|}{2\pi},&\text{when $|\alpha|>2\pi(n-1)$}.\end{array}\right.$
It’s worth noticing that the Toledo number of a DT representation is always contained in the interval $(-1,1)\setminus\{0\}$ and can be arbitrarily close to 00, highlighting another strong contrast with Fuchsian representations.
The topology of relative $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ character varieties has been studied by Mondello. Among other things, he characterized all their compact components, proving that the only compact connected components that aren’t isolated points are DT components.
Corollary 2.11 ([Mon16, Corollary 4.17]).
The relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains at most one compact connected component and it does precisely in the following cases:
  • $|\alpha|\in 2\pi\operatorname{\mathbb{Z}}$ and the compact component is an isolated point corresponding to an orthogonal representation.
  • $|\alpha|\in(0,2\pi)\cup(2\pi(n-1),2\pi n)$ and the compact component is the DT component from Theorem 2.9.
In the context of punctured spheres, Theorem A can now be reformulated as follows. If an element $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a reduced totally elliptic representation, then $\alpha$ is one of the angle vectors from Corollary 2.11 and $[\rho]$ belongs to the unique compact connected component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. Actually, Deroin–Tholozan already observed that totally elliptic representations belong to a bounded region of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
Proposition 2.12 ([DT19, Proposition 2.5]).
The closure of the subset of all conjugacy classes of totally elliptic representations inside $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is compact.
Deroin–Tholozan’s argument to prove Proposition 2.12 uses that all the representations in a sequence of totally elliptic representations are dominated by the same Fuchsian representation and therefore correspond to a sequence of Lipschitz equivariant maps $\operatorname{\mathbb{H}}\to\operatorname{\mathbb{H}}$ by a result of Guéritaud–Kassel [GK17]. The conclusion then follows from the Arzelà–Ascoli theorem (see [DT19] for more details).
Even though totally elliptic representations are contained in a bounded region, it could nevertheless be possible a priori to have a compact subset of a non-compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ made of totally elliptic representations. To finish the proof of Theorem A, we shall prove that no such compact subsets exist. We’ll do so by going over all possible values of $n$ inductively (recall that $n$ is the number of punctures of $\Sigma$).
2.3.1. The case $n=3$
A $3$-punctured sphere $\Sigma$ only contains three non-trivial simple closed curves: the three peripheral curves. So any point in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is automatically the conjugacy class of a totally elliptic representation. It turns out that when $n=3$, the relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either empty or a singleton. More precisely, let’s pick a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3}$ satisfying $c_{1}c_{2}c_{3}=1$. Now, if $\rho$ is a reduced totally elliptic representation, then the unique fixed points of $\rho(c_{1})$, $\rho(c_{2})$, and $\rho(c_{3})$ form a triangle in $\operatorname{\mathbb{H}}$ with vertices $(C_{1},C_{2},C_{3})$. Denote the rotation angles of $\rho(c_{1})$, $\rho(c_{2})$, and $\rho(c_{3})$ by $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})\in(0,2\pi)^{3}$. When we study the location of the points $(C_{1},C_{2},C_{3})$, we observe that the triangle they form can only be in three possible configurations, as explained in [Mar24, Section 3.1 and Table 1].
  • $|\alpha|\in\{2\pi,4\pi\}$ and the triangle is degenerate with $C_{1}=C_{2}=C_{3}$.
  • $|\alpha|<2\pi$ and the triangle $(C_{1},C_{2},C_{3})$ is anti-clockwise oriented with interior angles given by $(\alpha_{1}/2,\alpha_{2}/2,\alpha_{3}/2)$.
  • $|\alpha|>4\pi$ and the triangle $(C_{1},C_{2},C_{3})$ is clockwise oriented with interior angles given by $(\pi-\alpha_{1}/2,\pi-\alpha_{2}/2,\pi-\alpha_{3}/2)$.
If $|\alpha|\in\{2\pi,4\pi\}$, then $\rho$ is orthogonal. When $|\alpha|<2\pi$ or when $|\alpha|>4\pi$, then $\rho$ is a DT representation. (Note that when $|\alpha|\in(2\pi,4\pi)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is empty.) So, Theorem A holds for 3-punctured spheres.
2.3.2. The case $n=4$
Throughout this section, $\Sigma$ will denote a $4$-punctured sphere. We start by considering the case of representations with a non-peripheral simple closed curve in their kernel.
Lemma 2.13.
If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a totally elliptic representation with a non-peripheral simple closed curve in its kernel, then $\rho$ is orthogonal.
Proof.
We work with a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3},c_{4}$ satisfying $c_{1}c_{2}c_{3}c_{4}=1$. Let’s assume that $c_{1}c_{2}$ is a fundamental group element that represents the non-peripheral simple closed curve in the kernel of $\rho$. This means that there exist two elements $X,Y$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ such that $\rho(c_{1})=X^{-1}$, $\rho(c_{2})=X$, $\rho(c_{3})=Y$, and $\rho(c_{4})=Y^{-1}$. Now, observe that the fundamental group element $c_{2}c_{3}c_{1}c_{3}^{-1}$ represents a simple closed curve on $\Sigma$ (it’s the image of the simple closed curve represented by $c_{1}c_{2}$ under the Dehn twist along $c_{2}c_{3}$). Since $\rho$ is totally elliptic, we conclude that $\rho(c_{2}c_{3}c_{1}c_{3}^{-1})=[X,Y]$ is elliptic. By Lemma 2.1, this is possible only if $[X,Y]=1$, proving that $\rho$ is orthogonal. ∎
It was already observed in [Mar22, Remark 2.8] that if $|\alpha|<2\pi$ or if $|\alpha|>6\pi$, then every totally elliptic representation in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is a DT representation. It turns out that a stronger statement is true: a representation that maps a single non-peripheral simple closed curve to a regular elliptic element is necessarily a DT representation. The proof of this statement uses the characterization of DT representations in terms of Toledo number (given by (2.3)) and the additive properties of the Toledo number (see [Mar22] for more details). When $|\alpha|\in[2\pi,6\pi]$, however, there exist representations outside of the compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ that map some non-peripheral simple closed curves to elliptic elements, but are not totally elliptic (an explicit example is provided by Palesi in [Pal14, Section 5.4]).
We now give a more general argument that applies for every value of $|\alpha|$ and proves that totally elliptic representations always belong to a compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$. Let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a totally elliptic representation and consider the pure mapping class group orbit of $[\rho]$ inside $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ (we refer the reader to [Mar22] for a detailed description of this action). The orbit of $[\rho]$ is a bounded subset of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by Proposition 2.12. Bounded orbits for 4-punctured spheres have been studied by Cantat–Loray in the more general context of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ representations. Their work has the following implication.
Corollary 2.14 ([CL09, Theorem C]).
Let $\Sigma$ be a 4-punctured sphere. The relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an infinite bounded mapping class group orbit if and only if $|\alpha|<2\pi$ or $|\alpha|>6\pi$. Moreover, each infinite bounded orbit is dense in the DT component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
This means that if the mapping class group orbit of $[\rho]$ is infinite, then $|\alpha|<2\pi$ or $|\alpha|>6\pi$, and $\rho$ is a DT representation. If instead the orbit of $[\rho]$ is finite, then we can consult Lisovyy–Tykhyy’s classification of finite mapping class group orbits for 4-punctured spheres from [LT14] and observe that $\rho$ is either a representation into $\operatorname{PSO}(2)$ (which corresponds to the isolated points of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ when $|\alpha|\in\{2\pi,4\pi,6\pi\}$) or a DT representation. Altogether, this proves Theorem A for $4$-punctured spheres.
2.3.3. The case $n\geq 5$
Let now $\Sigma$ denote a sphere with $n\geq 5$ punctures and $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced totally elliptic representation. We’ll see that the proof of Theorem A in this case can be brought back to the case of 4-punctured spheres (Section 2.3.2). As before, we start with the case of representations with a non-peripheral simple closed curve in their kernel.
Lemma 2.15.
If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a reduced totally elliptic representation with a non-peripheral simple closed curve in its kernel, then $\rho$ is orthogonal.
Proof.
We choose a geometric generating family of $\pi_{1}\Sigma$ consisting of generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$. They can be chosen in a way that the simple closed curve in the kernel of $\rho$ is represented by the fundamental group element $c_{1}\cdots c_{i+1}$ for some $i\in\{2,\ldots,n-3\}$. This implies that $\rho(c_{i+1})=\rho(c_{1}\cdots c_{i})^{-1}$ and $\rho(c_{i+2})=\rho(c_{i+3}\cdots c_{n})^{-1}$. Arguing as in the proof of Lemma 2.13, we know that we can conjugate $\rho$ so that $\rho(c_{i+1})$ and $\rho(c_{i+2})$ both belong to $\operatorname{PSO}(2)$.
Now, note that the fundamental group element $c_{i+1}c_{j}(c_{1}\cdots c_{i})c_{j}^{-1}$ represents a simple closed curve for every $j\geq i+2$ and similarly for $c_{i+2}^{-1}c_{j}^{-1}(c_{i+3}\cdots c_{n})^{-1}c_{j}$ when $j\leq i+1$. The first one is the image of $c_{1}\cdots c_{i+1}$ by the Dehn twist along $c_{i+1}c_{j}$ and the second one is the image of $(c_{i+2}\cdots c_{n})^{-1}$ by the Dehn twist along $c_{j}c_{i+2}$. Since $\rho$ is totally elliptic, we conclude that $\rho(c_{i+1}c_{j}(c_{1}\cdots c_{i})c_{j}^{-1})=[\rho(c_{i+1}),\rho(c_{j})]$ and $\rho(c_{i+2}^{-1}c_{j}^{-1}(c_{i+3}\cdots c_{n})^{-1}c_{j})=[\rho(c_{i+2})^{-1},\rho(c_{j})^{-1}]$ are both elliptic, hence trivial by Lemma 2.1. Because we’re assuming $\rho$ to be reduced, it holds that $\rho(c_{i+1})\neq 1$ and $\rho(c_{i+2})\neq 1$. We conclude that $\rho(c_{j})$ belongs to $\operatorname{PSO}(2)$ for every $j$ again by Lemma 2.1. This proves that $\rho$ is orthogonal. ∎
From now on, let’s assume that $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel ($\rho$ is “regularly totally elliptic” in the terminology introduced after Definition 2.10). To such a representation corresponds a chain of triangles in the hyperbolic plane $\operatorname{\mathbb{H}}$. It’s constructed from a chained pants decomposition $\mathcal{B}$ of $\Sigma$, by which we mean a pants decomposition of $\Sigma$ in which each pair of pants contains at least one puncture. We then choose a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$, such that the $n-3$ pants curves of $\mathcal{B}$ are represented by the fundamental group elements $b_{i}=(c_{1}\cdots c_{i+1})^{-1}$ for $i=1,\ldots,n-3$. This is always possible as explained in [Mar24, Appendix B]. Since $\rho$ is reduced and has no simple closed curves in its kernel, all the elliptic elements $\rho(c_{1}),\ldots,\rho(c_{n})$ and $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are regular, and have a unique fixed point in $\operatorname{\mathbb{H}}$ which we denote by $C_{1},\ldots,C_{n}$ and $B_{1},\ldots,B_{n-3}$. The triangle chain of $\rho$ consists of the $n-2$ triangles
$(C_{1},C_{2},B_{1}),(B_{1},C_{3},B_{2}),\ldots,(B_{n-4},C_{n-2},B_{n-3}),(B_{n-3},C_{n-1},C_{n}).$
We call it the $\mathcal{B}$-triangle chain of $\rho$. For instance, when $n=6$, the $\mathcal{B}$-triangle chain of $\rho$ may look like this.
[…]
For more details on triangle chains, the reader is referred to [Mar24].
Every triangle in the chain of $\rho$ is associated to one of the $n-2$ pairs of pants determined by $\mathcal{B}$ in the same way as we described in Section 2.3.1. So, each triangle can be either degenerate to a single point, or clockwise or anti-clockwise oriented. It turns out that being a DT representation can be characterized by the orientations of the triangles.
Proposition 2.16 ([Mar24, Lemma 3.5]).
If there exists a chained pants decomposition $\mathcal{B}$ such that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate and have the same orientation, then $\rho$ is a DT representation. Conversely, if $\rho$ is a DT representation, then for every chained pants decomposition $\mathcal{B}$, all non-degenerate triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation.
If the $\mathcal{B}$-triangle chain of $\rho$ only contains degenerate triangles, then all $\rho(c_{i})$ fix the same point. This means that $\rho$ is an orthogonal representation. Instead, if at least one triangle is non-degenerate, then the elliptic elements $\rho(c_{i})$ are not all rotations around the same point in $\operatorname{\mathbb{H}}$ and $\rho$ is not orthogonal. In the latter case, we can actually replace $\mathcal{B}$ by a new chained pants decomposition of $\Sigma$ and find a compatible geometric presentation of $\pi_{1}\Sigma$ such that the new triangle chain of $\rho$ only consists of non-degenerate triangles. This new pants decomposition can be constructed using the same algorithm as in [FM23, Proposition 2] and the compatible geometric presentation of $\pi_{1}\Sigma$ is obtained by the same reasoning as in [FM23, Proposition 3]. From now on, we’ll therefore assume that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate.
We’ll denote the rotation angles of the images of the pants curve $b_{1},\ldots,b_{n-3}$ in $\mathcal{B}$ under $\rho$ by $\beta_{1},\ldots,\beta_{n-3}\in(0,2\pi)$. Recall that $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are all regular elliptic elements because we’re assuming that the kernel of $\rho$ doesn’t contain any simple closed curve.
Proposition 2.17.
If $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel, then all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation. Hence, $\rho$ is a DT representation by Proposition 2.16.
Proof.
For each $i=1,\ldots,n-3$, we consider the 4-punctured spheres $\Sigma^{(i)}\subset\Sigma$ with peripheral curves $((b_{i-1})^{-1},c_{i+1},c_{i+2},b_{i+1})$.
$b_{i-1}$$b_{i}$$b_{i+1}$$c_{i}$$c_{i+1}$$c_{i+2}$$c_{i+3}$
The sphere $\Sigma$ admits $n-3$ such sub-spheres, one for each pants curve in $\mathcal{B}$. We follow the convention of writing $b_{0}=c_{1}^{-1}$ and $\beta_{0}=2\pi-\alpha_{1}$, as well as $b_{n-2}=c_{n}$ and $\beta_{n-2}=\alpha_{n}$. The conjugacy class of the restriction of $\rho$ to $\Sigma^{(i)}$ is a point in the relative character variety
$\operatorname{Rep}_{\alpha^{(i)}}(\Sigma^{(i)},\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$
where $\alpha^{(i)}=(2\pi-\beta_{i-1},\alpha_{i+1},\alpha_{i+2},\beta_{i+1})$. The restriction of $\rho$ to $\Sigma^{(i)}$ is obviously still totally elliptic.
The standard pants decomposition of $\Sigma^{(i)}$ associated to the geometric system of generators $((b_{i-1})^{-1},c_{i+1},c_{i+2},b_{i+1})$ of $\pi_{1}\Sigma^{(i)}$ consists of one curve: $b_{i}$. We’ll write it $\mathcal{B}_{i}$. The $\mathcal{B}_{i}$-triangle chain of the restriction of $\rho$ to $\Sigma^{(i)}$ is made of the two triangles that share the vertex $B_{i}$ in the $\mathcal{B}$-triangle chain of $\rho$. These two triangles are non-degenerate because we’re assuming that every triangle in the $\mathcal{B}$-triangle chain of $\rho$ is non-degenerate. In particular, this means that none of the restrictions of $\rho$ to any of the spheres $\Sigma^{(i)}$ is a representation into $\operatorname{PSO}(2)$. So, by the discussion of Section 2.3.2 about 4-punctured spheres, they’re all DT representations because they are totally elliptic. This means that the two triangles in the $\mathcal{B}_{i}$-triangle chain of $\rho$ have the same orientation by Proposition 2.16. Now, because the $\mathcal{B}$-triangle chain of $\rho$ is made of the $\mathcal{B}_{i}$-triangle chain of each restriction, with an overlap of one non-degenerate triangle for consecutive indices, we conclude that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation. ∎
Proposition 2.17 finishes the proof of Theorem A.

### v2

@@
 Abstract.
-We characterize all totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ by showing that they are either representations into a compact subgroup or Deroin–Tholozan representations.
+A surface group representation into a Lie group is called totally elliptic if every simple closed curve on the surface is mapped to an elliptic element of the target group.
+In this note, we characterize all totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ and $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ by showing that they are either representations into a compact subgroup or Deroin–Tholozan representations.
 1. Introduction
-1.1. Main result
+1.1. Main results
 Surface group representations are homomorphisms from the fundamental group of a surface $\Sigma$ into a Lie group $G$.
-Their deformation spaces are called character varieties.
-A character variety is not necessarily a connected space.
-Dynamical behaviour and qualitative properties of representations may vary considerably between components.
-At one end of the spectrum are the higher Teichmüller spaces which are a generalization of the classical and extensively studied Teichmüller space of $\Sigma$.
-Higher Teichmüller spaces, as defined in [Wie18], consist of discrete and faithful representations and tend to behave like moduli spaces of geometric structures on $\Sigma$, although a precise correspondence is still unavailable.
+The topological quotient of the space of all homomorphisms $\pi_{1}\Sigma\to G$ by the conjugation action of $G$ will be called the character variety of the pair $(\Sigma,G)$ (even though the quotient may not be a Hausdorff topological space) and denoted by
+$\operatorname{Rep}(\Sigma,G)=\operatorname{Hom}(\pi_{1}\Sigma,G)/G.$
+A character variety is not necessarily a connected space; dynamical behaviour and qualitative properties of representations may vary considerably between components.
+At one end of the spectrum are components of discrete and faithful representations (so-called higher Teichmüller spaces [Wie18]).
 This article aims at drawing the reader’s attention towards the other end of the spectrum where connected components are mostly made of non-discrete representations.
-We’ll focus on a particular prototype of representations that are antipodal to discrete and faithful ones: the so-called totally elliptic representations.
+We will focus on specific prototype of surface group representations—known as totally elliptic representations—that are antipodal to discrete and faithful ones.
 Definition 1.1.
 A surface group representation $\pi_{1}\Sigma\to G$ is totally elliptic if every simple closed curve on $\Sigma$ is mapped to an elliptic element of $G$.
-By an elliptic element of $G$, we mean an element of a compact subgroup of $G$ [^1] Those elements are sometimes called compact in the literature.
-The surface $\Sigma$ will always be taken to be oriented and connected, and can be punctured.
-Recall that a simple closed curve on $\Sigma$ is the free isotopy class of a closed loop on $\Sigma$ that doesn’t self-intersect and isn’t homotopically trivial.
-A simple closed curve can be homotopic to a puncture, in which case we call it peripheral.
-By definition, the images of every peripheral curve by a totally elliptic representation are elliptic.
-We want to avoid the situation where peripheral curves are mapped to the identity and we say that a representation is reduced when this isn’t the case (Definition 2.7).
-The first instances of totally elliptic representations, though not the most interesting, are those into a compact subgroup of $G$.
-Interestingly, however, by requiring only simple closed curves to be mapped to elliptic elements (as opposed to all curves), one can construct Zariski dense examples of totally elliptic representations into non-compact Lie groups, such as every Hermitian Lie group (Section 1.3), whose deformations remain totally elliptic.
-The initial case to investigate is when $G$ is $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
-In that case, the first totally elliptic representations with Zariski dense image were discovered by Benedetto–Goldman [BG99], and later generalized by Deroin–Tholozan [DT19].
-Those representations were originally named supra-maximal in [DT19], but we’ll refer to them as DT representations (Definition 2.10), to avoid any confusion with maximal representations from higher Teichmüller theory.
-DT representations only exist when $\Sigma$ is a genus-0 surface with at least three punctures.
-As Deroin–Tholozan observed, their deformation spaces are compact components inside the corresponding relative character varieties and made exclusively of totally elliptic representations (Theorem 2.9).
-The question was raised whether they’re the only totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, besides representations into compact subgroups.
+By an elliptic element of $G$, we mean an element of a compact subgroup of $G$ (sometimes called a compact element of $G$).
+The surface $\Sigma$ will always be assumed to be oriented and connected, and may be punctured.
+A simple closed curve on $\Sigma$ is the free isotopy class of a closed loop on $\Sigma$ that does not self-intersect and is homotopically non-trivial.
+A simple closed curve can be homotopic to a puncture, in which case we call it a peripheral curve.
+By definition of a totally elliptic representation, the images of all peripheral curves are elliptic.
+We want to avoid the situation where peripheral curves are mapped to the identity element in $G$ and we say that a representation is reduced when this is not the case (Definition 2.7).
+The first examples of totally elliptic representations, though not the most striking, are those whose image lies in a compact subgroup of $G$.
+Interestingly, however, by requiring only simple closed curves to be mapped to elliptic elements (as opposed to all curves), one can construct totally elliptic representations with dense image into non-compact Lie groups and whose deformations remain totally elliptic.
+This turns out to be possible whenever $G$ is a Hermitian Lie group for instance (Section 1.4).
+This paper focuses on the case where $G$ is $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ or $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$.
+The first totally elliptic representations with (Zariski) dense image in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ were discovered by Benedetto–Goldman [BG99] for 4-punctured spheres, and later generalized by Deroin–Tholozan [DT19] for spheres with arbitrarily many punctures.
+Those representations were originally named supra-maximal in [DT19], but we shall call them as DT representations (Definition 2.10), to avoid any confusion with maximal representations from higher Teichmüller theory.
+Deroin–Tholozan proved that deforming conjugacy classes of DT representations, while preserving the rotation angle (Section 2.2) of the peripheral monodromies, leads to compact components inside the corresponding relative character varieties (Section 2.3) which are made exclusively of totally elliptic representations (Theorem 2.9).
+The question was raised whether DT representations are the only totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, besides representations into compact subgroups which we call orthogonal representations.
 We bring a positive answer to that question.
 Theorem A.
-Let $\Sigma$ be a connected and oriented surface with genus $g\geq 0$ and $n\geq 0$ punctures.
-Reduced totally elliptic representations $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ can only be of two kinds:
-  • The image of $\rho$ is contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. In that case, we say that $\rho$ is an orthogonal representation.
-  • If $\rho$ is not orthogonal, then $g=0$, $n\geq 3$, and $\rho$ is a DT representation.
+Consider a surface $\Sigma$ of genus $g\geq 0$ with $n\geq 0$ punctures, and let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced totally elliptic representation.
+If $\rho$ is not orthogonal, then $\Sigma$ must be a sphere with at least three punctures ($g=0$, $n\geq 3$), and $\rho$ is necessarily a DT representation.
 We prove Theorem A in Section 2.
+The proof can be sketched as follows.
+First, we notice that surfaces with genus at least $1$ and different than the closed torus admit simple closed curves that can be represented by the commutator of two fundamental group elements, each representing a simple closed curve (Lemma 2.2).
+However, the commutator of two elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is elliptic if and only if it is trivial (Lemma 2.1).
+These two observations will obstruct the existence of non-orthogonal totally elliptic representations when $g\geq 1$.
+The more challenging case is when $g=0$.
+When $\Sigma$ is a punctured sphere, we will prove that totally elliptic and non-orthogonal representations are DT representations by inducting on the number $n$ of punctures on $\Sigma$.
+The base case $n=4$ is a consequence of a combination of results by Deroin–Tholozan (Proposition 2.12) and Cantat–Loray (Corollary 2.15).
+For the induction step, we will require the notion of triangle chains developed in [Mar24] and the associated characterization of DT representations (Proposition 2.17).
+1.2. Applications and generalizations
+An immediate consequence of Theorem A is that DT representations can be characterized inside their relative character variety by the property of being totally elliptic.
+Corollary B.
+If $\Sigma$ is a sphere with $n\geq 3$ punctures and $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ maps every peripheral curve on $\Sigma$ to a non-trivial elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, then
+[…]
+The direct implication in Corollary B was established in [DT19, Lemma 3.2] (see also [Mar22, Proposition 2.7]) and the return implication follows immediately from Theorem A.
+The statement of Corollary B can be refined depending on the peripheral data of $\rho$.
+The angle of rotation of the image under $\rho$ of a peripheral curve on $\Sigma$ is a called a peripheral angle (Section 2.3).
+A peripheral angle is always a number in $(0,2\pi)$.
+When the sum of all peripheral angles of $\rho$ (one for each puncture of $\Sigma$) is not an integer multiple of $2\pi$, then $\rho$ cannot be orthogonal (see Corollary 2.11 for more details).
+In that case, the statement of Corollary B can be refined to:
+If the sum of the peripheral angles of $\rho$ is not an integer multiple of $2\pi$, then
+$\rho\text{ is a DT representation}\quad\Longleftrightarrow\quad\rho\text{ is totally elliptic.}$.
+In Section 3, we will generalize the statement of Theorem A to surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$—the complexification of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+Maximal compact subgroups of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ are conjugate to the projective unitary group $\operatorname{PSU}(2)$ and surface group representations into a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ are accordingly called unitary representations.
+In order to generalize Theorem A from $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ to $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$, we need to distinguish between irreducible and reducible representations.
+Recall that a representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ is reducible if its image is conjugate to a subgroup made of upper triangular elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$; it is irreducible otherwise.
+Despite the unfortunate linguistic resemblance, the notion of reducible representations and that of reduced representations (Definition 2.7) are different.
+Theorem C (Propositions 3.13.2).
+Consider a surface $\Sigma$ of genus $g\geq 0$ and with $n\geq 0$ punctures, and let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ be a reduced totally elliptic representation.
+  • If $\rho$ is irreducible, then either $\rho$ is unitary or $\Sigma$ must be sphere with at least three punctures ($g=0$, $n\geq 3$), and $\rho$ is conjugate to the composition of a DT representation with the inclusion $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}\subset\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$.
+  • If $\rho$ is reducible and $g\geq 1$, then the image of $\rho$ is conjugate to a subgroup made of diagonal elements in $\operatorname{PSU}(2)$.
+Theorem C leaves open the case of reducible totally elliptic representations into $\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ for genus-0 surfaces.
+As it turns out, such representations are not necessarily unitary.
+Theorem D (Example 3.4).
+If $\Sigma$ is a sphere with at least three punctures ($g=0$, $n\geq 3$), then there exist reduced totally elliptic representations $\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ that are reducible but not unitary.
 1.3. Motivations
-Characterizing a certain type of surface group representations (such as totally elliptic ones) by the images of simple closed curves may seem baroque at first.
+Characterizing a certain type of surface group representations by the images of simple closed curves, as we did in Corollary B, may seem baroque at first.
 Similar problems, however, have sparked substantial interest in recent decades.
-Bowditch famously raised the question of whether discrete and faithful representations of a closed surface group $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—also known as Fuchsian representations—can be characterized by the property of mapping every simple closed curve on $\Sigma$ to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (an element whose trace is larger than $2$ in absolute value) [Bow98, Question C].
-If true, this would be a finer characterization of Fuchsian representations which are known to be the only representations that map every non-trivial element of $\pi_{1}\Sigma$ to hyperbolic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
-Bowditch’s question has been answered positively for surfaces of genus $2$ by Marché–Wolff [MW16, MW19] and remains open in higher genera.
+Bowditch famously raised the question of whether discrete and faithful representations of a closed surface group into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—also known as Fuchsian representations—can be characterized by the property of mapping every simple closed curve to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (an element whose trace is larger than two in absolute value).
+In analogy to Definition 1.1, such representations will be called totally hyperbolic.
+Question 1 ([Bow98, Question C]).
+Let $\Sigma$ denote a closed surface of genus at least two. Is the following equivalence true for a representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$:
+$\rho\text{ is a Fuchsian representation}\quad\Longleftrightarrow\quad\rho\text{ is totally hyperbolic}\quad?$
+Bowditch’s question (Question 1) has been answered positively for surfaces of genus two by Marché–Wolff [MW16, MW19] and remains open in higher genera.
+A positive answer to Bowditch’s question would be a finer characterization of Fuchsian representations which are already known to be exactly those representations mapping every non-trivial element of $\pi_{1}\Sigma$ to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+A surface group representation into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ whose image consists of hyperbolic elements and the identity is sometimes called purely hyperbolic, so that Fuchsian representations corresponds to faithful purely hyperbolic representations.
 Of course, not every representation is Fuchsian.
 Goldman proved that the character variety of representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma$ is a closed surface of genus $g\geq 2$, has $34g-$ 3 connected components [Gol88].
-Two of the components are made of conjugacy classes of Fuchsian representations, forming two copies of the Teichmüller space of $\Sigma$; the other $54g-$ 5 components are known as intermediate components and they don’t contain any Fuchsian representation.
-Bowditch’s question is equivalent to asking whether every representation whose conjugacy class lies in an intermediate component maps at least one simple closed curve on $\Sigma$ to a non-hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+The components are distinguished by an integer-valued invariant called Euler number.
+The two components with extremal Euler number are made of conjugacy classes of Fuchsian representations and form two copies of the Teichmüller space of $\Sigma$; the other $54g-$ 5 components are known as intermediate components and they do not contain any Fuchsian representation.
+Bowditch’s question (Question 1) is equivalent to asking whether every representation whose conjugacy class lies in an intermediate component maps at least one simple closed curve on $\Sigma$ to a non-hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
 Marché–Wolff also proved that a positive answer to Bowditch’s question would imply that the mapping class group action on every intermediate component is ergodic [MW16, Theorem 1.6].
-The latter statement is known as Goldman’s ergodicity conjecture [Gol06, Conjecture 3.1] and is open, like Bowditch’s question, for surfaces of genus at least $3$.
-Goldman also conjectured that almost every non-Fuchsian closed surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ should be the holonomy of a branched hyperbolic structure on $\Sigma$ which is the data of hyperbolic metric of $\Sigma$ with several conical singularities of angle $2\pi k$ with $k\geq 2$.
-To be precise, according to Goldman, every dense representation whose conjugacy class lies in an intermediate component should be the holonomy of a branched hyperbolic structure [Gol06, Conjecture 3.9].
-Mathews proved some partial implication from a positive answer to Bowditch’s question to Goldman’s geometrization conjecture for intermediate components of co-Euler class $1$ [Mat12].
-Observe as well that Goldman’s ergodicity conjecture implies a “full measure version” of the geometrization conjecture, as well as a positive answer to Bowditch’s question on a full measure subset of intermediate components.
+The latter statement is known as Goldman’s ergodicity conjecture [Gol06, Conjecture 3.1] and is open, like Bowditch’s question, for surfaces of genus at least three.
+Goldman also conjectured that almost every non-Fuchsian closed surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ should be the holonomy of a branched hyperbolic structure on $\Sigma$ which is the data of a hyperbolic metric of $\Sigma$ with several conical singularities of angle $2\pi k$ with $k\geq 2$.
+To be precise, according to Goldman, every dense representation whose conjugacy class lies in an intermediate component with non-zero Euler number should be the holonomy of a branched hyperbolic structure [Gol06, Conjecture 3.9].
+Mathews proved some partial implication from a positive answer to Bowditch’s question to Goldman’s geometrization conjecture for the intermediate component of Euler number $32g-$ 3 [Mat12] and Faraco covered the case of purely hyperbolic representations [Far21].
+Bowditch’s question (Question 1) and Goldman’s conjectures on intermediate components are intertwined.
+In addition to Marché–Wolff’s results, we point out that Goldman’s ergodicity conjecture implies a “full measure version” of the geometrization conjecture, as well as a positive answer to Bowditch’s question on a full measure subset of intermediate components.
 Bowditch’s question Geometrization conjecture Ergodicity conjecture [MW16]
-Goldman’s two conjectures and Bowditch’s question, and the interplay among them, constitute a whole program to study intermediate components which we refer to as the Bowditch–Goldman program.
-Totally elliptic representations, and more particularly DT representations, naturally arise as a prototype of surface group representations that fit perfectly in the Bowditch–Goldman program.
+Goldman’s two conjectures and Bowditch’s question, as well as the interplay among them, constitute a whole program to study intermediate components which we refer to as the Bowditch–Goldman program.
+In this context, totally elliptic representations—particularly DT representations—align naturally with the broader philosophy of the Bowditch–Goldman program, even though, strictly speaking, the program concerns closed surfaces.
 By definition, they map every simple closed curve to an elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, and therefore satisfy the analogue of Bowditch’s condition on intermediate representations in a strong sense.
@@
 As explained in [DT19, Section 4] and further detailed in [FM23], DT representations are also holonomies of hyperbolic cone structures with branch point singularities, thereby satisfying a precise version of the geometrization conjecture.
-1.3. Perspectives away from $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$
-This note is also an opportunity to survey some recent developments about totally elliptic representations into Lie groups $G$ different than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
-Several families of compact connected components of totally elliptic representations were identified when $G$ is a Hermitian Lie group.
-They’re all genus-0 surface group representations.
+The study of totally elliptic surface group representations is also motivated by the recent focus on finite mapping class group orbits in character varieties (see for instance [LLL23, BM24, LT14] and references therein) which led to a complete understanding of finite orbits for surface group representations into $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$.
+Although not all finite orbits of conjugacy classes of surface group representations into $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ are made of totally elliptic representations, most of them are (see [BM24] for more details).
+Totally elliptic representations can thus be viewed as a natural generalization of representations whose conjugacy classes lie in finite mapping class group orbits.
+1.4. Perspectives away from $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$
+This note also provides an opportunity to survey recent developments concerning totally elliptic representations into Lie groups $G$ other than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+Several families of connected components in relative character varieties—all consisting entirely of conjugacy classes of totally elliptic representations—have been identified in the case where $G$ is a Hermitian Lie group.
+These components are all compact and arise specifically for punctured spheres.
 The first examples were found by Tholozan–Toulisse when $G$ is the Lie group $\mathrm{SU}(p,q)$, as well as when $G$ is $\mathrm{Sp}(2n,\operatorname{\mathbb{R}})$ or $\mathrm{SO}^{\ast}(2n)$ [TT21].
 Analogous compact components have been identified by Feng–Zhang when $G$ is the identity component of $\mathrm{SO}(2,n)$ [FZ23].
-Their methods studies the topology of some moduli spaces of parabolic Higgs bundles which relate to relative character varieties via the non-abelian Hodge correspondence.
-In particular, they proved that for every representation $\rho\colon\pi_{1}\Sigma\to G$ whose conjugacy class belongs to one of the compact components they identified, and for every punctured Riemann sphere structure on $\Sigma$, there exists a $\rho$-equivariant holomorphic map from the universal cover $\widetilde{\Sigma}$ to the symmetric space $G/K$.
-The same property for DT representations was proven by Mondello [Mon16].
-Surface group representations satisfying this condition are sometimes called universal variations of Hodge structures.
-Unlike DT representations which are all Zariski dense, it’s not clear yet whether every compact component of totally elliptic representations into Hermitian Lie groups other than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ contains a Zariski dense representation (almost all of them do though, by the preponderance of Zariski dense representations in any character variety).
-The topology of those compact components is also yet to be determined, whereas DT components are known to be homeomorphic (even symplectomorphic) to complex projective spaces [DT19, Theorem 4].
-There are however more explicit examples of compact components of totally elliptic representations discovered by Marc-Zwecker [MZ24].
-Marc-Zwecker’s examples are made of representations of a pair of pants into $\mathrm{SU}(2,1)$ where the peripheral curves are mapped to products of reflections through three complex geodesic lines in the complex hyperbolic space of real dimension four.
-The three complex geodesic lines are the mirrors and they’re arranged as a triangle.
+Both group of authors studied the topology of relative character varieties via the non-abelian Hodge correspondence and identified compact components in this way.
+Not only did they prove that every representation $\rho\colon\pi_{1}\Sigma\to G$ whose conjugacy class belongs to one of the compact components they found is totally elliptic, but they also proved that the local system associated to $\rho$ supports a universal variation of Hodge structure.
+This means that for every punctured Riemann sphere structure on $\Sigma$, there exists a $\rho$-equivariant holomorphic map from the universal cover $\widetilde{\Sigma}$ to the symmetric space $G/K$.
+The same property for DT representations was established by Mondello [Mon16].
+In contrast to DT representations, which all have Zariski dense image, it remains unclear whether every compact component described in [TT21, FZ23] contains a Zariski dense representation (though almost all of them do, given the prevalence of Zariski dense representations in character varieties).
+Moreover, the topology of these compact components is still not fully understood, whereas DT components are known to be homeomorphic, and even symplectomorphic, to complex projective spaces [DT19, Theorem 4].
+There are however more explicit examples of components in relative character varieties made of totally elliptic representations that were discovered by Marc-Zwecker [MZ24].
+Marc-Zwecker’s components are also compact and made of representations of a pair of pants into $\mathrm{SU}(2,1)$ where the peripheral curves are mapped to products of reflections through three complex geodesic lines in the complex hyperbolic space of real dimension four.
+The three complex geodesic lines are the mirrors and they are arranged as a triangle.
 The resulting components are homeomorphic to 2-dimensional spheres.
-The geometric nature of the representations in Marc-Zwecker’s components may lead to other compact components for spheres with a higher number of punctures obtained by “chaining” the triangles of mirrors in the spirit of [Mar24].
-A natural question to ask is whether the class of Lie groups for which there exist substantial connected components entirely made of totally elliptic genus-0 surface group representations is larger than the class of Hermitian Lie groups.
-Question 1.
-Are there semisimple real Lie groups other than Hermitian Lie groups for which there exist compact connected components (of maximal expected dimension) inside relative character varieties of genus-0 surface group representations made of totally elliptic representations and containing Zariski dense representations?
-There are reasons to believe that the key property is the presence of a nonempty open subset of elliptic elements in the target Lie group.
-This property is shared by all Hermitian Lie groups, but not only; for instance, Sugiura proved that $\mathrm{SO}(2p,q)$ has an open subset of elliptic elements [Sug59, Theorem 8], even though it’s not Hermitian if $p\geq 2$ and $q\geq 3$.
-In contrast, the subset of elliptic elements in $\mathrm{SL}_{n}\operatorname{\mathbb{R}}$ has empty interior for every $n\geq 3$.
-Furthermore, Kabenyuk proved that admitting a nonempty open subset of elliptic elements is equivalent to having a compact Cartan subgroup [Kab90].
-One should nevertheless be aware that “large” Lie groups, unlike $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$, may admit subgroups $\Gamma$ containing only elliptic elements but such that $\Gamma$ is not contained into a compact subgroup.
+The geometric nature of the representations in Marc-Zwecker’s components may lead to other components of totally elliptic representations for spheres with a higher number of punctures obtained by “chaining” the triangles of mirrors in the spirit of [Mar24].
+A natural question to ask is whether the class of Lie groups for which there exist (non-trivial) connected components entirely made of totally elliptic surface group representations is larger than the class of Hermitian Lie groups.
+It is nevertheless unclear whether such components are necessarily compact when the target group is not $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (compare Proposition 2.12), but it seems likely to.
+Question 2.
+Are there semisimple real Lie groups other than Hermitian Lie groups for which there exist compact connected components of maximal expected dimension in relative character varieties which are made of totally elliptic representations and contain Zariski dense representations?
+One should nevertheless be aware that “large” Lie groups, unlike $\mathrm{SL}_{2}\operatorname{\mathbb{R}}$ or $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$, may admit a subgroup $\Gamma$ containing only elliptic elements—a purely elliptic subgroup—but such that $\Gamma$ is not contained in a compact subgroup.
+[^1] A proof that every purely elliptic subgroup of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ is conjugate to a subgroup of $\mathrm{SU}(2)$ can be found in [JS87, Theorem 2.13.1].
+The analogous statement for $\mathrm{SL}_{2}\operatorname{\mathbb{R}}$ and $\mathrm{SO}(2)$ is a direct consequence of Lemma 2.1.
 Bass gives an example of such a $\Gamma$ which is isomorphic to a rank-2 free subgroup of $\mathrm{SL}_{3}\operatorname{\mathbb{C}}$ [Bas80, Counterexample 1.10].
-Bass’ example is not Zariski dense and, in fact, it couldn’t be as explained for instance in [Pra94].
+Bass’ example is not Zariski dense and, in fact, it could not be as explained for instance in [Pra94].
 However, when $\operatorname{\mathbb{C}}$ is replaced by a non-Archimedean valued field $\mathbb{F}$, then the situation is similar to the case of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ as Parreau proved that a finitely generated subgroup of $\mathrm{SL}_{n}\mathbb{F}$ containing only elliptic elements is necessarily bounded [Par03, Théorème 1].
-This doesn’t rule out the existence of interesting totally elliptic surface group representations into $\mathrm{SL}_{n}\mathbb{F}$.
-1.4. Acknowledgments
+This, of course, does not rule out the existence of totally elliptic surface group representations into $\mathrm{SL}_{n}\mathbb{F}$ whose image is not bounded.
+1.5. Acknowledgements
 Many thanks to Samuel Bronstein and Nicolas Tholozan for enriching conversations on totally elliptic representations.
@@
 I also extend my gratitude to my friends Jacques Audibert and Xenia Flamm for their insightful comments on an earlier draft of this paper and for pointing out to me Parreau’s paper.
-2. Proof of Theorem A
+2. Totally elliptic representations in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$
 2.1. Overview
-The proof of Theorem A unfolds over the next subsections.
+This section contains the proof of Theorem A which unfolds over the next subsections.
 We start by showing that when the genus of $\Sigma$ is at least $1$, then the only totally elliptic representations are orthogonal representations (Section 2.2).
-The arguments mostly rely on the observation that a commutator of elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is itself elliptic only when it’s trivial (Lemma 2.1).
-We then briefly recall what DT representations are (Section 2.3) and prove that a totally elliptic genus-0 surface group representation that isn’t orthogonal must be a DT representation (Sections 2.3.12.3.3).
+The arguments mostly rely on the observation that a commutator of elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is itself elliptic only when it is trivial (Lemma 2.1).
+We then briefly recall what DT representations are (Section 2.3) and prove that a totally elliptic genus-0 surface group representation that is not orthogonal must be a DT representation.
+We start with the case of spheres with three or four punctures (Sections 2.3.1 and 2.3.2), and then proceed by induction to prove the statement for spheres with an arbitrary number of punctures (Section 2.3.3).
 2.2. Positive genus
-We start by proving that a totally elliptic representation $\rho\colon\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of a surface $\Sigma$ of genus $g\geq 1$ is necessarily an orthogonal representation (Propositions 2.5 and 2.8).
+We start by proving that a totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of a surface $\Sigma$ of genus $g\geq 1$ is necessarily an orthogonal representation (Propositions 2.5 and 2.8).
 Most of the reasoning will rely on the following standard results about $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we include for the sake of completeness.
 According to Definition 1.1, the elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ fall into two categories: the regular elliptic elements and the identity.
-The regular elliptic elements are those whose trace is smaller than $2$ in absolute value.
+The regular elliptic elements are those whose trace is smaller than two in absolute value.
 Geometrically, they act on the hyperbolic plane $\operatorname{\mathbb{H}}$ by rotations around their unique fixed point.
-The subset of regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is homeomorphic to the open ball $(0,2\pi)\times\operatorname{\mathbb{H}}$, where the first factor records the rotation angle of a regular elliptic and the second its unique fixed point (the centre of rotation).
-Lemma 2.1.
-A regular elliptic element $A$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is contained in a unique maximal compact subgroup which consists of all the elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ which fix the unique fixed point of $A$.
-Furthermore, if $X$ is any element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, then the commutator
+The subset of regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is homeomorphic to the open ball $(0,2\pi)\times\operatorname{\mathbb{H}}$, where the first factor is the rotation angle and the second is the fixed point (the centre of rotation).
+A regular elliptic element $A\in\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is contained in a unique maximal compact subgroup which is a conjugate of $\mathrm{PSO}(2)$.
+It is therefore abelian and consists of all the elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ which fix the unique fixed point of $A$.
+It also coincides with the centralizer of $A$ in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+Lemma 2.1 (see e.g. [Bea83, Theorem 7.39.2]).
+If $A\in\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is regular elliptic and $X\in\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is any element, then the commutator
 $[A,X]=AXA^{-1}X^{-1}$
-is either the identity or a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. The commutator is trivial if and only if $X$ is elliptic and fixes the unique fixed point of $A$.
+is either trivial or hyperbolic. Furthermore, the commutator is trivial if and only if $X$ is elliptic and fixes the unique fixed point of $A$.
 Proof.
-The unique maximal compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ that contains $A$ is its centralizer, which is conjugate to $\operatorname{PSO}(2)$.
-Since $A$ is regular, any element that commutes with $A$ is either the identity or a regular elliptic with the same fixed point as $A$.
-Let’s now consider the commutator $[A,X]$.
-As we just said, if $[A,X]=1$, then $X$ is elliptic and fixes the unique fixed point of $A$.
-If $[A,X]\neq 1$ and $o$ denotes the unique fixed point of $A$, then $XA^{-1}X^{-1}$ is regular elliptic with fixed point $X.o\neq o$.
+First, if $[A,X]=1$, then $X$ belongs to the centralizer of $A$, which means that $X$ is elliptic and fixes the unique fixed point of $A$.
+Now, if $[A,X]\neq 1$ and $o$ denotes the unique fixed point of $A$, then $XA^{-1}X^{-1}$ is regular elliptic with fixed point $X.o\neq o$.
 Note that if $\alpha$ is the rotation angle of $A$, then the rotation angle of $XA^{-1}X^{-1}$ is $2\pi-\alpha$.
@@
 Lemma 2.1 plays a central role in our discussion because it tells us that the commutator of two elliptic elements is elliptic if and only if it is trivial.
-It turns out that the commutator of two simple closed curves (seen as elements of $\pi_{1}\Sigma$) sometimes represents a simple closed curve too.
-We must, however, be careful, because a simple closed curve on $\Sigma$ is represented by many elements in $\pi_{1}\Sigma$, all conjugate to each other.
+It turns out that two simple closed curves with geometric intersection number one can be lifted to specific fundamental group elements whose commutator represents a simple closed curve too.
+We must however be careful, because a simple closed curve on $\Sigma$ is represented by many elements in $\pi_{1}\Sigma$, all conjugate to each other.
 Lemma 2.2.
-If two simple closed curves on $\Sigma$ intersect once (their geometric intersection number is 1) and we represent them in the simplest way by two fundamental group elements $a$ and $b$ with basepoint being the intersection point, then $[a,b]$ also represents a simple closed curve on $\Sigma$.
+Let $\alpha$ and $\beta$ be two simple closed loops on $\Sigma$ intersecting once at $x\in\Sigma$ and such that the geometric intersection number of their free homotopy classes is one.
+We denote by $a$ and $b$ the fundamental group elements with basepoint $x$ associated to $\alpha$ and $\beta$.
+If $[a,b]$ is a non-trivial fundamental group element (which is always the case if $\Sigma$ is not a closed torus), then it represents a simple closed curve on $\Sigma$.
 Proof.
-Pick two loops on $\Sigma$ that only intersect at the basepoint and whose homotopy classes are $a$ and $b$.
-Enlarge each loop by replacing it with a thin ribbon.
-The boundary of the union of the two ribbons is an embedded circle in $\Sigma$ which is in the free homotopy class determined by the fundamental group element $[a,b]$.
+Enlarge the two loops $\alpha$ and $\beta$ by replacing them with a thin ribbon.
+The union of the two ribbons is homeomorphic to a torus with one boundary component which is an embedded circle in $\Sigma$.
+The free homotopy class of that circle is non-trivial as long as $\Sigma$ is not a closed torus, and it corresponds to the free homotopy class determined by the fundamental group element $[a,b]$.
 ∎
-As we’re about to see, Lemmas 2.1 and 2.2 together are the main obstruction for constructing totally elliptic surface group representations.
-Let’s start by introducing some notation.
+As we are about to see, Lemmas 2.1 and 2.2 together provide the main obstruction for constructing totally elliptic surface group representations.
+Let us start by introducing some notation.
 We pick a geometric generating family for $\pi_{1}\Sigma$ which consists in $2g+n$ generators $a_{1},\ldots,a_{g}$, $b_{1},\ldots,b_{g}$, and $c_{1},\ldots,c_{n}$ satisfying the unique relation
 $\prod_{i=1}^{g}[a_{i},b_{i}]=\prod_{i=1}^{n}c_{i}.$
-A possible choice of geometric generators are the homotopy classes of the loops illustrated on Figure 1.
+A possible choice of geometric generators is given by the homotopy classes of the loops illustrated on Figure 1.
+[…]
 Figure 1. A system of geometric generators for a surface $\Sigma$ of genus $g$ with $n$ punctures.
@@
 More precisely, if $\gamma_{1}$ and $\gamma_{2}$ are two loops on $\Sigma$ based at the same point, then their concatenation $\gamma_{1}\gamma_{2}$ is the loop obtained by going along $\gamma_{2}$ first, and then along $\gamma_{1}$.
-This rule applies to fundamental group multiplications too.
+This rule applies for fundamental group multiplications too.
 Lemma 2.4.
-If the genus $g$ of $\Sigma$ is at least 2 and $i\neq j$ are two indices in $\{1,\ldots,g\}$, then the following fundamental group elements represent simple closed curves on $\Sigma$:
+If $\Sigma$ is a surface of genus $g\geq 2$ and $i\neq j$ are two distinct indices in $\{1,\ldots,g\}$, then the following fundamental group elements represent simple closed curves on $\Sigma$:
 $[a_{i},b_{i}],\quad[a_{j},b_{j}],\quad[b_{i}^{-1}b_{j}^{-1},a_{i}],\quad[a_{i}^{-1}a_{j}^{-1},b_{j}],$
@@
 Proof.
-Each pair of fundamental group elements appearing in a given commutator represents two simple closed curves on $\Sigma$ with geometric intersection number $1$.
-So, by Lemma 2.2, we can expect their commutators to represent simple closed curves too.
-We must, however, be careful when we represent a simple closed curve by a fundamental group element because of basepoint issues.
-The choice we made ensures that each commutator indeed represents a simple closed curve on $\Sigma$.
-This can be seen by drawing the corresponding curves on Figure 1.
-∎
+Each pair of fundamental group elements appearing in one of the commutators above represents two simple closed curves on $\Sigma$ with geometric intersection number one.
+By Lemma 2.2, we can thus expect their commutators to represent simple closed curves too.
+This is however not automatic because of the issues we mentioned before stating Lemma 2.2.
+Thankfully, the way we picked the generators $a_{1},\ldots,a_{g}$ and $b_{1},\ldots,b_{g}$ ensures that each commutator does indeed represent a simple closed curve.
+This can be seen by drawing the corresponding loops using the generators of Figure 1 and the concatenation conventions described in Remark 2.3, like in the figure below.
+[Uncaptioned $i$$j$$[b_{i}^{-1}b_{j}^{-1},a_{1}]$
 Proposition 2.5.
-Assume that $\Sigma$ is closed and has genus $g\geq 1$.
+Assume that $\Sigma$ is a closed surface of genus $g\geq 1$.
 If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a totally elliptic representation, then $\rho$ is an orthogonal representation.
 Proof.
-We’ll write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$.
+We shall write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$.
 If $g=1$, then the image of $\rho$ is generated by $A_{1}$ and $B_{1}$ which satisfy $[A_{1},B_{1}]=1$.
 So, the image of $\rho$ is generated by two commuting elliptic elements, hence contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
-Let’s now consider the case $g\geq 2$.
-If $\rho$ is the trivial representation, then it’s orthogonal.
+Let us now consider the case $g\geq 2$.
+If $\rho$ is the trivial representation, then it is orthogonal.
 If not, then there exists an index $i\in\{1,\ldots,g\}$ such that either $A_{i}\neq 1$ or $B_{i}\neq 1$.
-We’ll look at the case where $A_{i}\neq 1$ (the case where $B_{i}\neq 1$ is identical).
+We will look at the case where $A_{i}\neq 1$ (the case where $B_{i}\neq 1$ is identical).
 Let $j\neq i$ be another index.
-Because $\rho$ is totally elliptic, it maps all the simple closed curves from Lemma 2.4 to the identity because of Lemma 2.1.
+Since $\rho$ is totally elliptic, it maps all the simple closed curves from Lemma 2.4 to the identity because of Lemma 2.1.
 In particular, we have $[A_{i},B_{i}]=1$ and $[A_{i},B_{i}^{-1}B_{j}^{-1}]=1$, implying that $[A_{i},B_{j}]=1$.
 Since $[A_{i},B_{i}]=1$, up to conjugating $\rho$, we may assume that $A_{i}$ and $B_{i}$ belong to $\operatorname{PSO}(2)$.
-From $[A_{i},B_{j}]=1$ and our assumption that $A_{i}\neq 1$, we conclude that $B_{j}\in\operatorname{PSO}(2)$ by Lemma 2.1.
+From $[A_{i},B_{j}]=1$ and our assumption that $A_{i}\neq 1$, we conclude that $B_{j}\in\operatorname{PSO}(2)$.
 We also infer from Lemmas 2.1 and 2.4 that $[A_{j},B_{j}]=[A_{j},B_{j}^{-1}B_{i}^{-1}]=[A_{j},B_{i}A_{i}^{-1}B_{j}]=1$.
-If $B_{j}\neq 1$, then we immediately conclude that $A_{j}\in\operatorname{PSO}(2)$ from $[A_{j},B_{j}]=1$ by Lemma 2.1.
+If $B_{j}\neq 1$, then we immediately conclude that $A_{j}\in\operatorname{PSO}(2)$ from $[A_{j},B_{j}]=1$.
 Since $[A_{j},B_{j}]=1$ and $[A_{j},B_{j}^{-1}B_{i}^{-1}]=1$, we deduce that $[A_{j},B_{i}]=1$ which also implies $A_{j}\in\operatorname{PSO}(2)$ as soon as $B_{i}\neq 1$.
@@
 A similar observation holds for the second fundamental group element in (2.1) which is equal to $b_{i}a_{i}b_{i}a_{i}^{-1}b_{i}^{-1}\cdot c_{j}^{-1}b_{i}^{-1}c_{j}$ and is thus a product of conjugates of $b_{i}$ and $b_{i}^{-1}$.
-By drawing a rapid picture, we convince ourselves that both represent simple closed curves on $\Sigma$.
-We prove in the same way that the fundamental group element (2.2) represents a simple closed curve too.
-∎
+We can draw closed loops that represent these two fundamental group elements, as well as the one in (2.2), using the generators of Figure 1 and the concatenation rules of Remark 2.3, like in the figure below.
+This shows that they all three represent simple closed curves on $\Sigma$.
+[Uncaptioned $i$$j$$a_{i}^{-1}[a_{i},b_{i}]c_{j}^{-1}a_{i}c_{j}$[Uncaptioned $i$$j$$k$$b_{i}c_{k}b_{i}^{-1}c_{j}b_{i}a_{i}^{-1}b_{i}^{-1}c_{k}^{-1}b_{i}^{-1}c_{j}^{-1}$
 Definition 2.7.
-A surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is said to be reduced if it doesn’t map any peripheral curve on $\Sigma$ to the identity.
-In terms of the geometric generators from Figure 1, a representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is reduced if it doesn’t map any generator $c_{i}$ to the identity.
-Note that every non-reduced totally elliptic surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ corresponds to a reduced totally elliptic representation $\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma^{\prime}$ is obtained from $\Sigma$ by removing the punctures whose peripheral curves are mapped to the identity.
+A surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ is said to be reduced if it does not map any peripheral curve on $\Sigma$ to the identity.
+Equivalently and in terms of the geometric generators from Figure 1, $\rho$ is reduced if it does not map any of the generators $c_{1},\ldots,c_{n}$ to the identity.
+Every non-reduced surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ factorizes through a reduced representation $\rho^{\prime}\colon\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$, where $\Sigma^{\prime}$ is obtained by filling in the punctures of $\Sigma$ whose corresponding peripheral curves are mapped to the identity by $\rho$.
+More precisely, the inclusion $\Sigma\subset\Sigma^{\prime}$ induces a surjective morphism $p\colon\pi_{1}\Sigma\twoheadrightarrow\pi_{1}\Sigma^{\prime}$ and $\rho^{\prime}\circ p=\rho$.
+Note that if $\rho$ is totally elliptic, then so is $\rho^{\prime}$.
 Proposition 2.8.
 Assume that $\Sigma$ is a surface of genus $g\geq 1$ and with $n\geq 1$ punctures.
-If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a reduced totally elliptic representation, then $\rho$ is an orthogonal representation.
+If $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is a reduced totally elliptic representation, then $\rho$ is orthogonal.
 Proof.
-As before, we’ll write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$, as well as $C_{j}=\rho(c_{j})$.
+As before, we shall write $A_{i}=\rho(a_{i})$ and $B_{i}=\rho(b_{i})$, as well as $C_{j}=\rho(c_{j})$.
 Proposition 2.5 implies that all the $A_{i}$ and $B_{i}$ are contained in the same orthogonal subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, which we can assume to be $\operatorname{PSO}(2)$ up to conjugating $\rho$.
 In particular, $[A_{i},B_{i}]=1$ for every index $i$.
-Let’s first assume that there exists an index $i\in\{1,\ldots,g\}$ such that $A_{i}\neq 1$ or $B_{i}\neq 1$.
-We’ll consider the case where $A_{i}\neq 1$ (similar arguments apply when $B_{i}\neq 1$).
-Using the simple closed curves (2.1) from Lemma 2.6 and since $\rho$ is totally elliptic, we conclude that $A_{i}^{-1}[A_{i},B_{i}]C_{j}^{-1}A_{i}C_{j}$ is elliptic for every index $j$.
+Let us first assume that there exists an index $i\in\{1,\ldots,g\}$ such that $A_{i}\neq 1$ or $B_{i}\neq 1$.
+We will consider the case where $A_{i}\neq 1$ (similar arguments apply when $B_{i}\neq 1$).
+Using the simple closed curves (2.1) from Lemma 2.6 and since $\rho$ is totally elliptic, we conclude that $A_{i}^{-1}[A_{i},B_{i}]C_{j}^{-1}A_{i}C_{j}$ is elliptic for every index $j\in\{1,\ldots,n\}$.
 Because $[A_{i},B_{i}]=1$, we even have that $[A_{i}^{-1},C_{j}^{-1}]$ is elliptic.
@@
 This holds for every index $j$ and therefore the image of $\rho$ is contained in $\operatorname{PSO}(2)$.
-Now, let’s consider the case where $A_{i}=B_{i}=1$ for every $i\in\{1,\ldots,g\}$.
-Since we’re assuming $\rho$ to be reduced, we must have $n\geq 2$.
-Using the simple closed curve (2.2) from Lemma 2.6, we deduce that $B_{i}C_{k}B_{i}^{-1}C_{j}B_{i}A_{i}^{-1}B_{i}^{-1}C_{k}^{-1}B_{i}^{-1}C_{j}^{-1}$ is elliptic for every index $i$ and $j>k$.
-Since $A_{i}=B_{i}=1$, we even have $[C_{k},C_{j}]=1$ for every $j>k$ in $\{1,\ldots,n\}$.
+Now, let us consider the case where $A_{i}=B_{i}=1$ for every $i\in\{1,\ldots,g\}$.
+Since we are assuming $\rho$ to be reduced, we must have $n\geq 2$.
+Using the simple closed curve (2.2) from Lemma 2.6, we deduce that $B_{i}C_{k}B_{i}^{-1}C_{j}B_{i}A_{i}^{-1}B_{i}^{-1}C_{k}^{-1}B_{i}^{-1}C_{j}^{-1}$ is elliptic for every index $i\in\{1,\ldots,g\}$ and every $j>k$ in $\{1,\ldots,n\}$.
+Since $A_{i}=B_{i}=1$, we obtain $[C_{k},C_{j}]=1$ for every $j>k$ in $\{1,\ldots,n\}$.
 This implies that $\rho$ is orthogonal for the same reasons as before.
 ∎
-Propositions 2.5 and 2.8 say that the only examples of reduced totally elliptic representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ when $\Sigma$ has positive genus and an arbitrary number of punctures are orthogonal representations.
+Propositions 2.5 and 2.8 say that when $\Sigma$ has positive genus and an arbitrary number of punctures, then every reduced totally elliptic representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is orthogonal.
+In order to finish the proof of Theorem A, we shall now prove that if the genus of $\Sigma$ is zero, then every reduced totally elliptic surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is either orthogonal or a DT representation.
 2.3. Genus zero
-In order to finish the proof of Theorem A, we shall prove that a reduced totally elliptic genus-0 surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is either orthogonal or a DT representation.
-Clearly, if $\Sigma$ has at most two punctures, then $\rho$ is orthogonal.
-Let’s now assume that $\Sigma$ is a sphere with a set $\mathcal{P}$ of $n\geq 3$ punctures.
-Recall that the simple closed curves on $\Sigma$ that are homotopic to a puncture are called peripheral.
-By definition, a reduced totally elliptic representation maps every peripheral curve on $\Sigma$ to a regular elliptic element.
-In other words, the conjugacy class $[\rho]$ of a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ belongs to an $\alpha$-relative character variety
-$\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$
-which consists of all conjugacy classes of representations mapping peripheral curves around every puncture $p\in\mathcal{P}$ to regular elliptic elements in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ of fixed rotation angle $\alpha_{p}\in(0,2\pi)$.
-The numbers $\alpha_{p}$ are the components of the angle vector $\alpha\in(0,2\pi)^{\mathcal{P}}$.
-We’ll write
+First, note that if $\Sigma$ has at most two punctures, then every totally elliptic surface group representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is orthogonal.
+So we will assume that $\Sigma$ is a sphere with a set $\mathcal{P}$ of $n\geq 3$ punctures.
+For every angle vector $\alpha\in(0,2\pi)^{\mathcal{P}}$, the subspace of the character variety $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ consisting of all conjugacy classes of representations mapping every peripheral curve around a puncture $p\in\mathcal{P}$ to regular elliptic elements of fixed rotation angle $\alpha_{p}\in(0,2\pi)$ is called $\alpha$-relative character variety and is denoted by
+$\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})\subset\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}}).$
+The numbers $\alpha_{p}$ are called peripheral angles and their sum is written
 $|\alpha|=\sum_{p\in\mathcal{P}}\alpha_{p}.$
-Unlike in the positive genus case, there are Zariski dense totally elliptic representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ when $\Sigma$ is an $n$-punctured sphere.
+By definition, a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ maps every peripheral curve on $\Sigma$ to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+The images under $\rho$ of all peripheral curves around a puncture $p\in\mathcal{P}$ are conjugate and thus have the same rotation angle $\alpha_{p}\in(0,2\pi)$.
+In other words, the conjugacy class $[\rho]$ of $\rho$ belongs to the $\alpha$-relative character variety where $\alpha$ is the vector of peripheral angles $\alpha_{p}$.
+In contrast to the positive genus case (see Propositions 2.5 and 2.8), there exist totally elliptic genus-0 surface group representations with Zariski dense image in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
 The first examples were discovered via computer simulations by Benedetto–Goldman in the case $n=4$.
-Their construction was later generalized by Deroin–Tholozan for spheres with an arbitrary number of punctures.
+Their construction was later generalized by Deroin–Tholozan for spheres with an arbitrary number of punctures, leading to the following combined statement.
 Theorem 2.9 ([BG99, DT19]).
-If $|\alpha|<2\pi$ or $|\alpha|>2\pi(n-1)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains a compact connected component isomorphic to $\mathbb{CP}^{n-3}$ and entirely made of conjugacy classes of Zariski dense totally elliptic representations that don’t contain any simple closed curve in their kernel.
+If $|\alpha|<2\pi$ or $|\alpha|>2\pi(n-1)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains a compact connected component isomorphic to $\mathbb{CP}^{n-3}$ and entirely made of conjugacy classes of totally elliptic representations with Zariski dense image.
+More precisely, these representations map every simple closed curves on $\Sigma$ to a regular elliptic element.
 Definition 2.10.
 A representation whose conjugacy class belongs to one of the compact components of Theorem 2.9 will be called a DT representation. The components themselves will be referred to as DT components.
-According to Theorem 2.9, DT representations are more than just totally elliptic in the sense of Definition 1.1; they are “regularly totally elliptic” by which we mean that every simple closed curve is mapped to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+According to the last statement of Theorem 2.9, DT representations are more than just totally elliptic in the sense of Definition 1.1; they are “regularly totally elliptic” by which we mean that every simple closed curve is mapped to a regular elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
 DT representations can be characterized by their Toledo number in the sense of Burger–Iozzi–Wienhard [BIW10].
-Namely, according to Deroin–Tholozan’s original definition, a conjugacy class of representations in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ belongs to a DT component if and only if its Toledo number is
+Namely, according to Deroin–Tholozan’s original definition, a point in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ belongs to a DT component if and only if its Toledo number is
 (2.3) $\left\{\begin{array}[]{ll}1-\frac{|\alpha|}{2\pi},&\text{when $|\alpha|<2\pi$},\\ (n-1)-\frac{|\alpha|}{2\pi},&\text{when $|\alpha|>2\pi(n-1)$}.\end{array}\right.$
-It’s worth noticing that the Toledo number of a DT representation is always contained in the interval $(-1,1)\setminus\{0\}$ and can be arbitrarily close to 00, highlighting another strong contrast with Fuchsian representations.
+It is worth noticing that the Toledo number of a DT representation is always contained in $(-1,1)\setminus\{0\}$ and can be arbitrarily close to 00 when $|\alpha|$ approaches $2\pi$ or $2\pi(n-1)$.
+Since the absolute value of the Toledo number on $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ can be as high as $-\chi(\Sigma)=n-2$ according to [BIW10], DT representations exhibit nearly minimal Toledo number among all points of $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
+This highlights another strong contrast with Fuchsian representations which have maximal Toledo (or Euler) number.
 The topology of relative $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ character varieties has been studied by Mondello.
-Among other things, he characterized all their compact components, proving that the only compact connected components that aren’t isolated points are DT components.
+Among other things, he characterized all their compact components, proving that the only compact connected components that are not isolated points are DT components.
 Corollary 2.11 ([Mon16, Corollary 4.17]).
-The relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains at most one compact connected component and it does precisely in the following cases:
+The $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains at most one compact connected component and it does precisely in the following cases:
   • $|\alpha|\in 2\pi\operatorname{\mathbb{Z}}$ and the compact component is an isolated point corresponding to an orthogonal representation.
-  • $|\alpha|\in(0,2\pi)\cup(2\pi(n-1),2\pi n)$ and the compact component is the DT component from Theorem 2.9.
-In the context of punctured spheres, Theorem A can now be reformulated as follows.
-If an element $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a reduced totally elliptic representation, then $\alpha$ is one of the angle vectors from Corollary 2.11 and $[\rho]$ belongs to the unique compact connected component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
-Actually, Deroin–Tholozan already observed that totally elliptic representations belong to a bounded region of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
+  • $|\alpha|\in(0,2\pi)\cup(2\pi(n-1),2\pi n)$ and the compact component is one of the DT components from Theorem 2.9.
+With Corollary 2.11 in mind and in order to conclude the proof of Theorem A, we shall show that if an element $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a reduced totally elliptic representation, then $\alpha$ is one of the angle vectors from Corollary 2.11 and $[\rho]$ belongs to the unique compact connected component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
+We will use a result of Deroin–Tholozan who observed that conjugacy classes of totally elliptic representations remain in a bounded region of the character variety.
 Proposition 2.12 ([DT19, Proposition 2.5]).
-The closure of the subset of all conjugacy classes of totally elliptic representations inside $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is compact.
-Deroin–Tholozan’s argument to prove Proposition 2.12 uses that all the representations in a sequence of totally elliptic representations are dominated by the same Fuchsian representation and therefore correspond to a sequence of Lipschitz equivariant maps $\operatorname{\mathbb{H}}\to\operatorname{\mathbb{H}}$ by a result of Guéritaud–Kassel [GK17].
-The conclusion then follows from the Arzelà–Ascoli theorem (see [DT19] for more details).
-Even though totally elliptic representations are contained in a bounded region, it could nevertheless be possible a priori to have a compact subset of a non-compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ made of totally elliptic representations.
-To finish the proof of Theorem A, we shall prove that no such compact subsets exist.
-We’ll do so by going over all possible values of $n$ inductively (recall that $n$ is the number of punctures of $\Sigma$).
+The subset of all conjugacy classes of totally elliptic representations inside the character variety $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is bounded (i.e.
+contained in a compact subset).
+Deroin–Tholozan’s argument to prove Proposition 2.12 goes as follows.
+First, consider the closure of the subset of totally elliptic representations inside $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ and call it $\overline{\mathcal{TE}}$.
+The set $\overline{\mathcal{TE}}$ is the collection of all representations in $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ that map every simple closed curve on $\Sigma$ to either an elliptic or a parabolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+All the representations in $\overline{\mathcal{TE}}$ are dominated (in the sense of [GKW15, DT16]) by every Fuchsian representation in $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by a result of Guéritaud–Kassel [GK17].
+So, if we fix an arbitrary Fuchsian representation $j\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, it will dominate every representation in a sequence $(\rho_{k})\subset\overline{\mathcal{TE}}$.
+Domination means that to every $\rho_{k}$ corresponds a $(j,\rho_{k})$-equivariant 1-Lipschitz map $\operatorname{\mathbb{H}}\to\operatorname{\mathbb{H}}$.
+The existence of a converging subseqeunce for $(\rho_{k})$ now follows from the Arzelà–Ascoli theorem (see [DT19] for more details).
+This argument shows that $\overline{\mathcal{TE}}\subset\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is sequentially compact, hence compact because $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is a metrizable space.
+Now, if $p$ denotes the quotient map from $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ to the non-Hausdorff quotient $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$, then the subset of all conjugacy classes of totally elliptic representations in $\operatorname{Rep}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is contained in the compact set $p(\overline{\mathcal{TE}})$.
+Even though totally elliptic representations remain in a bounded region by Proposition 2.12, it could nevertheless be possible that some non-compact components of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contain conjugacy classes of totally elliptic representations.
+To conclude the proof of Theorem A, we shall prove that this is not the case.
+We will do so using an induction on the number $n$ of punctures of $\Sigma$.
 2.3.1. The case $n=3$
-A $3$-punctured sphere $\Sigma$ only contains three non-trivial simple closed curves: the three peripheral curves.
-So any point in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is automatically the conjugacy class of a totally elliptic representation.
-It turns out that when $n=3$, the relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either empty or a singleton.
-More precisely, let’s pick a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3}$ satisfying $c_{1}c_{2}c_{3}=1$.
-Now, if $\rho$ is a reduced totally elliptic representation, then the unique fixed points of $\rho(c_{1})$, $\rho(c_{2})$, and $\rho(c_{3})$ form a triangle in $\operatorname{\mathbb{H}}$ with vertices $(C_{1},C_{2},C_{3})$.
-Denote the rotation angles of $\rho(c_{1})$, $\rho(c_{2})$, and $\rho(c_{3})$ by $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})\in(0,2\pi)^{3}$.
+A $3$-punctured sphere $\Sigma$ only contains three non-trivial simple closed curves: its three peripheral curves.
+Any point in an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is thus automatically the conjugacy class of a totally elliptic representation.
+It turns out that when $n=3$, an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either empty or a singleton.
+More precisely, let us pick a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3}$ satisfying $c_{1}c_{2}c_{3}=1$, similarly as on Figure 1.
+Now, if $\rho$ is a reduced totally elliptic representation, then the unique fixed points of $\rho(c_{1})$, $\rho(c_{2})$, $\rho(c_{3})$ form a triangle in the hyperbolic plane with vertices $(C_{1},C_{2},C_{3})$.
+Denote the rotation angles of $\rho(c_{1})$, $\rho(c_{2})$, $\rho(c_{3})$ by $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})\in(0,2\pi)^{3}$.
 When we study the location of the points $(C_{1},C_{2},C_{3})$, we observe that the triangle they form can only be in three possible configurations, as explained in [Mar24, Section 3.1 and Table 1].
@@
 When $|\alpha|<2\pi$ or when $|\alpha|>4\pi$, then $\rho$ is a DT representation.
-(Note that when $|\alpha|\in(2\pi,4\pi)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is empty.) So, Theorem A holds for 3-punctured spheres.
+Finally, if $|\alpha|\in(2\pi,4\pi)$, then $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is empty.
+Altogether, this proves Theorem A for 3-punctured spheres.
 2.3.2. The case $n=4$
-Throughout this section, $\Sigma$ will denote a $4$-punctured sphere. We start by considering the case of representations with a non-peripheral simple closed curve in their kernel.
+Throughout this section, $\Sigma$ will denote a sphere with four punctures. We start by considering the case of representations with a non-peripheral simple closed curve in their kernel.
 Lemma 2.13.
@@
 We work with a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},c_{2},c_{3},c_{4}$ satisfying $c_{1}c_{2}c_{3}c_{4}=1$.
-Let’s assume that $c_{1}c_{2}$ is a fundamental group element that represents the non-peripheral simple closed curve in the kernel of $\rho$.
+Up to cyclically permuting the generators, we may assume that $c_{1}c_{2}$ is a fundamental group element that represents the non-peripheral simple closed curve in the kernel of $\rho$.
 This means that there exist two elements $X,Y$ of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ such that $\rho(c_{1})=X^{-1}$, $\rho(c_{2})=X$, $\rho(c_{3})=Y$, and $\rho(c_{4})=Y^{-1}$.
-Now, observe that the fundamental group element $c_{2}c_{3}c_{1}c_{3}^{-1}$ represents a simple closed curve on $\Sigma$ (it’s the image of the simple closed curve represented by $c_{1}c_{2}$ under the Dehn twist along $c_{2}c_{3}$).
-Since $\rho$ is totally elliptic, we conclude that $\rho(c_{2}c_{3}c_{1}c_{3}^{-1})=[X,Y]$ is elliptic.
-By Lemma 2.1, this is possible only if $[X,Y]=1$, proving that $\rho$ is orthogonal.
+We now claim that the fundamental group element $c_{1}c_{3}^{-1}c_{2}c_{3}$ represents a simple closed curve on $\Sigma$.
+This can be seen directly on Figure 1 by applying the concatenation conventions of Remark 2.3.
+Alternatively, apply a Dehn twist along the simple closed curve represented by $c_{2}c_{3}$ to the simple closed curve represented by $c_{1}c_{2}$.
+The resulting closed curve is simple and represented by $c_{1}c_{3}^{-1}c_{2}c_{3}$.
+Finally, using that $\rho$ is totally elliptic and that $c_{1}c_{3}^{-1}c_{2}c_{3}$ represents a simple closed curve, we conclude that $\rho(c_{1}c_{3}^{-1}c_{2}c_{3})=[X^{-1},Y^{-1}]$ is elliptic.
+By Lemma 2.1, this is possible only if $[X^{-1},Y^{-1}]=1$, proving that $\rho$ is orthogonal.
 ∎
-It was already observed in [Mar22, Remark 2.8] that if $|\alpha|<2\pi$ or if $|\alpha|>6\pi$, then every totally elliptic representation in $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is a DT representation.
-It turns out that a stronger statement is true: a representation that maps a single non-peripheral simple closed curve to a regular elliptic element is necessarily a DT representation.
-The proof of this statement uses the characterization of DT representations in terms of Toledo number (given by (2.3)) and the additive properties of the Toledo number (see [Mar22] for more details).
-When $|\alpha|\in[2\pi,6\pi]$, however, there exist representations outside of the compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ that map some non-peripheral simple closed curves to elliptic elements, but are not totally elliptic (an explicit example is provided by Palesi in [Pal14, Section 5.4]).
-We now give a more general argument that applies for every value of $|\alpha|$ and proves that totally elliptic representations always belong to a compact component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
-Let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a totally elliptic representation and consider the pure mapping class group orbit of $[\rho]$ inside $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ (we refer the reader to [Mar22] for a detailed description of this action).
-The orbit of $[\rho]$ is a bounded subset of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by Proposition 2.12.
-Bounded orbits for 4-punctured spheres have been studied by Cantat–Loray in the more general context of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ representations.
-Their work has the following implication.
+Having proved Lemma 2.13, we will now consider reduced totally elliptic representations $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ with no simple closed curve in their kernel.
+In the terminology introduced after Definition 2.10, they are regularly totally elliptic.
+In particular, each conjugacy class $[\rho]$ belongs to an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ for some angle vector $\alpha\in(0,2\pi)^{4}$.
+We want to prove that every such $\rho$ is either a DT or an orthogonal representation.
+Let us start with the case $|\alpha|<2\pi$ or $|\alpha|>6\pi$ which was already treated in [Mar22].
+It turns that mapping a single non-peripheral closed curve to an elliptic element is enough to be a DT representation in that case.
+Lemma 2.14 ([Mar22, Remark 2.8]).
+Let $\alpha\in(0,2\pi)^{4}$ be an angle vector such that $|\alpha|<2\pi$ or $|\alpha|>6\pi$ and let $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
+If $\rho$ maps any non-peripheral simple closed curve on $\Sigma$ to a regular elliptic element, then $\rho$ is a DT representation.
+In particular, every totally elliptic representation is a DT representation.
+The proof of Lemma 2.14 given in [Mar22] uses the characterization of DT representations in terms of Toledo number (given by (2.3)) and the additive properties of the Toledo number.
+When $|\alpha|\in[2\pi,6\pi]$, the situation is not as simple as in Lemma 2.14, because there exist representations that map some non-peripheral simple closed curves on $\Sigma$ to elliptic elements, but which are not totally elliptic.
+An explicit example is provided by Palesi in [Pal14, Section 5.4].
+We will nevertheless prove that if $|\alpha|\in[2\pi,6\pi]$ and $[\rho]\in\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is the conjugacy class of a totally elliptic representation, then actually $|\alpha|\in\{2\pi,4\pi,6\pi\}$ and $\rho$ is orthogonal.
+We will use an argument of dynamical flavour that turns out to also cover the case of Lemma 2.14.
+The pure mapping class group of $\Sigma$ acts on $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ for any angle vector $\alpha\in(0,2\pi)^{4}$ (details can be found in [Mar22]).
+If $\rho$ is totally elliptic, then every point in the pure mapping class group orbit of $[\rho]$ is the conjugacy class of a totally elliptic representation.
+The orbit of $[\rho]$ is therefore bounded by Proposition 2.12.
+There are now two possibilities.
+First, the orbit of $[\rho]$ may be finite.
+It turns out that finite orbits have been classified by Lisovyy–Tykhyy [LT14].
+Their classification shows that any finite orbit in an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ is either an isolated point or belongs to a DT component.
+Recall from Corollary 2.11 that $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an isolated point if and only if $|\alpha|\in\{2\pi,4\pi,6\pi\}$, and the isolated point is always the conjugacy class of an orthogonal representation.
+Secondly, we shall consider the case where the orbit of $[\rho]$ is an infinite bounded set.
+Infinite bounded orbits, in the context of 4-punctured spheres, have been studied by Cantat–Loray.
+A consequence of their work is the following statement.
+[^2] There is a small subtlety here.
+Cantat–Loray study the pure mapping class group action on the real points of the complex GIT quotient of $\operatorname{Hom}(\pi_{1}\Sigma,\mathrm{SL}_{2}\operatorname{\mathbb{C}})$ by $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ and not on the topological quotient of $\operatorname{Hom}(\pi_{1}\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ by $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ like we do.
+However, since all the representations whose conjugacy class lies in an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ are reductive, Corollary 2.15 is indeed a consequence of Cantat–Loray’s results.
 Corollary 2.15 ([CL09, Theorem C]).
-Let $\Sigma$ be a 4-punctured sphere.
-The relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an infinite bounded mapping class group orbit if and only if $|\alpha|<2\pi$ or $|\alpha|>6\pi$.
-Moreover, each infinite bounded orbit is dense in the DT component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
-This means that if the mapping class group orbit of $[\rho]$ is infinite, then $|\alpha|<2\pi$ or $|\alpha|>6\pi$, and $\rho$ is a DT representation.
-If instead the orbit of $[\rho]$ is finite, then we can consult Lisovyy–Tykhyy’s classification of finite mapping class group orbits for 4-punctured spheres from [LT14] and observe that $\rho$ is either a representation into $\operatorname{PSO}(2)$ (which corresponds to the isolated points of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ when $|\alpha|\in\{2\pi,4\pi,6\pi\}$) or a DT representation.
-Altogether, this proves Theorem A for $4$-punctured spheres.
+If an $\alpha$-relative character variety $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$ contains an infinite bounded orbit, then $|\alpha|<2\pi$ or $|\alpha|>6\pi$ and the orbit is dense in the DT component of $\operatorname{Rep}_{\alpha}(\Sigma,\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$.
+Consequently, no matter if the orbit of $[\rho]$ is finite or infinite, $\rho$ is always either a DT or an orthogonal representation. This proves Theorem A for 4-punctured spheres.
 2.3.3. The case $n\geq 5$
 Let now $\Sigma$ denote a sphere with $n\geq 5$ punctures and $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced totally elliptic representation.
-We’ll see that the proof of Theorem A in this case can be brought back to the case of 4-punctured spheres (Section 2.3.2).
 As before, we start with the case of representations with a non-peripheral simple closed curve in their kernel.
@@
 Proof.
-We choose a geometric generating family of $\pi_{1}\Sigma$ consisting of generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$.
-They can be chosen in a way that the simple closed curve in the kernel of $\rho$ is represented by the fundamental group element $c_{1}\cdots c_{i+1}$ for some $i\in\{2,\ldots,n-3\}$.
+We choose a geometric generating family of $\pi_{1}\Sigma$ consisting of generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$, similarly as on Figure 1.
+They can be chosen in a way that the simple closed curve in the kernel of $\rho$ is represented by the fundamental group element $c_{1}\cdots c_{i+1}$ for some $i\in\{1,\ldots,n-3\}$.
 This implies that $\rho(c_{i+1})=\rho(c_{1}\cdots c_{i})^{-1}$ and $\rho(c_{i+2})=\rho(c_{i+3}\cdots c_{n})^{-1}$.
-Arguing as in the proof of Lemma 2.13, we know that we can conjugate $\rho$ so that $\rho(c_{i+1})$ and $\rho(c_{i+2})$ both belong to $\operatorname{PSO}(2)$.
-Now, note that the fundamental group element $c_{i+1}c_{j}(c_{1}\cdots c_{i})c_{j}^{-1}$ represents a simple closed curve for every $j\geq i+2$ and similarly for $c_{i+2}^{-1}c_{j}^{-1}(c_{i+3}\cdots c_{n})^{-1}c_{j}$ when $j\leq i+1$.
-The first one is the image of $c_{1}\cdots c_{i+1}$ by the Dehn twist along $c_{i+1}c_{j}$ and the second one is the image of $(c_{i+2}\cdots c_{n})^{-1}$ by the Dehn twist along $c_{j}c_{i+2}$.
-Since $\rho$ is totally elliptic, we conclude that $\rho(c_{i+1}c_{j}(c_{1}\cdots c_{i})c_{j}^{-1})=[\rho(c_{i+1}),\rho(c_{j})]$ and $\rho(c_{i+2}^{-1}c_{j}^{-1}(c_{i+3}\cdots c_{n})^{-1}c_{j})=[\rho(c_{i+2})^{-1},\rho(c_{j})^{-1}]$ are both elliptic, hence trivial by Lemma 2.1.
-Because we’re assuming $\rho$ to be reduced, it holds that $\rho(c_{i+1})\neq 1$ and $\rho(c_{i+2})\neq 1$.
-We conclude that $\rho(c_{j})$ belongs to $\operatorname{PSO}(2)$ for every $j$ again by Lemma 2.1.
+We will argue as in the proof of Lemma 2.13.
+First, note that the fundamental group element $(c_{1}\cdots c_{i})c_{i+2}^{-1}c_{i+1}c_{i+2}$ represents a simple closed curve on $\Sigma$ that is the image of the simple closed curve represented by $c_{1}\cdots c_{i+1}$ under the Dehn twist along $c_{i+1}c_{i+2}$.
+Since we are assuming $\rho$ to be totally elliptic, it implies that $\rho\big{(}(c_{1}\cdots c_{i})c_{i+2}^{-1}c_{i+1}c_{i+2}\big{)}=\big{[}\rho(c_{i+1}^{-1}),\rho(c_{i+2}^{-1})\big{]}$ is elliptic.
+We deduce that $\big{[}\rho(c_{i+1}),\rho(c_{i+2})\big{]}=1$ from Lemma 2.1.
+So, up to conjugating $\rho$, we may assume that $\rho(c_{i+1})$ and $\rho(c_{i+2})$ both belong to $\operatorname{PSO}(2)$.
+Now, observe further that the fundamental group element $(c_{1}\cdots c_{i})c_{j}^{-1}c_{i+1}c_{j}$ represents a simple closed curve for every $j\geq i+2$ which is the image of the simple closed curve represented by $c_{1}\cdots c_{i+1}$ under the Dehn twist along $c_{i+1}c_{j}$.
+Similarly, for every $j\leq i+1$, $(c_{i+3}\cdots c_{n})^{-1}c_{j}c_{i+2}^{-1}c_{j}^{-1}$ represents the simple closed curve obtained by applying the Dehn twist along $c_{j}c_{i+2}$ to the simple closed curve represented by $(c_{i+2}\cdots c_{n})^{-1}$.
+Since $\rho$ is totally elliptic, we conclude that $\rho\big{(}(c_{1}\cdots c_{i})c_{j}^{-1}c_{i+1}c_{j}\big{)}=\big{[}\rho(c_{i+1})^{-1},\rho(c_{j})^{-1}\big{]}$ and $\rho\big{(}(c_{i+3}\cdots c_{n})^{-1}c_{j}c_{i+2}^{-1}c_{j}^{-1}\big{)}=\big{[}\rho(c_{i+2}),\rho(c_{j})\big{]}$ are both elliptic, hence trivial by Lemma 2.1.
+Because we are assuming $\rho$ to be reduced, it holds that $\rho(c_{i+1})\neq 1$ and $\rho(c_{i+2})\neq 1$.
+We conclude that $\rho(c_{j})$ belongs to $\operatorname{PSO}(2)$ for every index $j$.
 This proves that $\rho$ is orthogonal.
 ∎
-From now on, let’s assume that $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel ($\rho$ is “regularly totally elliptic” in the terminology introduced after Definition 2.10).
+From now on, we will assume that $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel.
+In the terminology introduced after Definition 2.10, $\rho$ is regularly totally elliptic.
 To such a representation corresponds a chain of triangles in the hyperbolic plane $\operatorname{\mathbb{H}}$.
-It’s constructed from a chained pants decomposition $\mathcal{B}$ of $\Sigma$, by which we mean a pants decomposition of $\Sigma$ in which each pair of pants contains at least one puncture.
+It is constructed from a chained pants decomposition $\mathcal{B}$ of $\Sigma$, by which we mean a pants decomposition of $\Sigma$ in which each pair of pants contains at least one puncture.
 We then choose a geometric presentation of $\pi_{1}\Sigma$ with generators $c_{1},\ldots,c_{n}$ satisfying $c_{1}\cdots c_{n}=1$, such that the $n-3$ pants curves of $\mathcal{B}$ are represented by the fundamental group elements $b_{i}=(c_{1}\cdots c_{i+1})^{-1}$ for $i=1,\ldots,n-3$.
@@
 Proposition 2.17 ([Mar24, Lemma 3.5]).
-If there exists a chained pants decomposition $\mathcal{B}$ such that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate and have the same orientation, then $\rho$ is a DT representation.
-Conversely, if $\rho$ is a DT representation, then for every chained pants decomposition $\mathcal{B}$, all non-degenerate triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation.
+Let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced regularly totally elliptic representation.
+If there exists a chained pants decomposition $\mathcal{B}$ of $\Sigma$ such that the $\mathcal{B}$-triangle chain of $\rho$ contains at least one non-degenerate triangle and all non-degenerate triangles have the same orientation, then $\rho$ is a DT representation.
+Conversely, if $\rho$ is a DT representation, then for every chained pants decomposition $\mathcal{B}$ of $\Sigma$, all non-degenerate triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation.
 If the $\mathcal{B}$-triangle chain of $\rho$ only contains degenerate triangles, then all $\rho(c_{i})$ fix the same point.
 This means that $\rho$ is an orthogonal representation.
-Instead, if at least one triangle is non-degenerate, then the elliptic elements $\rho(c_{i})$ are not all rotations around the same point in $\operatorname{\mathbb{H}}$ and $\rho$ is not orthogonal.
+Instead, if at least one triangle is non-degenerate, then the elliptic elements $\rho(c_{i})$ are not all rotations around the same point and $\rho$ is not orthogonal.
 In the latter case, we can actually replace $\mathcal{B}$ by a new chained pants decomposition of $\Sigma$ and find a compatible geometric presentation of $\pi_{1}\Sigma$ such that the new triangle chain of $\rho$ only consists of non-degenerate triangles.
 This new pants decomposition can be constructed using the same algorithm as in [FM23, Proposition 2] and the compatible geometric presentation of $\pi_{1}\Sigma$ is obtained by the same reasoning as in [FM23, Proposition 3].
-From now on, we’ll therefore assume that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate.
-We’ll denote the rotation angles of the images of the pants curve $b_{1},\ldots,b_{n-3}$ in $\mathcal{B}$ under $\rho$ by $\beta_{1},\ldots,\beta_{n-3}\in(0,2\pi)$.
-Recall that $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are all regular elliptic elements because we’re assuming that the kernel of $\rho$ doesn’t contain any simple closed curve.
+From now on, we will therefore assume that all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ are non-degenerate.
+We denote the rotation angles of the images of the pants curve $b_{1},\ldots,b_{n-3}$ in $\mathcal{B}$ under $\rho$ by $\beta_{1},\ldots,\beta_{n-3}\in(0,2\pi)$.
+Recall that $\rho(b_{1}),\ldots,\rho(b_{n-3})$ are all regular elliptic because we are assuming $\rho$ to be regularly totally elliptic.
 Proposition 2.18.
-If $\rho$ is a reduced totally elliptic representation with no simple closed curves in its kernel, then all the triangles in the $\mathcal{B}$-triangle chain of $\rho$ have the same orientation.
+Under our assumptions that $\rho$ is a reduced regularly totally elliptic representation and that its $\mathcal{B}$-triangle chain is made of non-degenerate triangles, then all triangles must have the same orientation.
 Hence, $\rho$ is a DT representation by Proposition 2.17.
@@
 The conjugacy class of the restriction of $\rho$ to $\Sigma^{(i)}$ is a point in the relative character variety
-$\operatorname{Rep}_{\alpha^{(i)}}(\Sigma^{(i)},\operatorname{PSL}_{2}\operatorname{\mathbb{R}})$
+$\operatorname{Rep}_{\alpha^{(i)}}(\Sigma^{(i)},\operatorname{PSL}_{2}\operatorname{\mathbb{R}}),$
 where $\alpha^{(i)}=(2\pi-\beta_{i-1},\alpha_{i+1},\alpha_{i+2},\beta_{i+1})$. The restriction of $\rho$ to $\Sigma^{(i)}$ is obviously still totally elliptic.
-The standard pants decomposition of $\Sigma^{(i)}$ associated to the geometric system of generators $((b_{i-1})^{-1},c_{i+1},c_{i+2},b_{i+1})$ of $\pi_{1}\Sigma^{(i)}$ consists of one curve: $b_{i}$.
-We’ll write it $\mathcal{B}_{i}$.
+The standard pants decomposition of $\Sigma^{(i)}$ associated to the geometric system of generators $((b_{i-1})^{-1},c_{i+1},c_{i+2},b_{i+1})$ of $\pi_{1}\Sigma^{(i)}$ consists of one simple closed curve represented by $b_{i}$.
+We will denote this pants decomposition by $\mathcal{B}_{i}$.
 The $\mathcal{B}_{i}$-triangle chain of the restriction of $\rho$ to $\Sigma^{(i)}$ is made of the two triangles that share the vertex $B_{i}$ in the $\mathcal{B}$-triangle chain of $\rho$.
-These two triangles are non-degenerate because we’re assuming that every triangle in the $\mathcal{B}$-triangle chain of $\rho$ is non-degenerate.
-In particular, this means that none of the restrictions of $\rho$ to any of the spheres $\Sigma^{(i)}$ is a representation into $\operatorname{PSO}(2)$.
-So, by the discussion of Section 2.3.2 about 4-punctured spheres, they’re all DT representations because they are totally elliptic.
+These two triangles are non-degenerate by assumption.
+In particular, this means that none of the restrictions of $\rho$ to any of the spheres $\Sigma^{(i)}$ is orthogonal.
+So, by the discussion of Section 2.3.2 about 4-punctured spheres, all the restrictions are DT representations because they are totally elliptic.
 This means that the two triangles in the $\mathcal{B}_{i}$-triangle chain of $\rho$ have the same orientation by Proposition 2.17.
@@
 Proposition 2.18 finishes the proof of Theorem A.
+3. Totally elliptic representations in $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$
+3.1. Overview
+The goal of this section is to classify totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$—the complexification of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—by proving Theorems C and D.
+We will first prove that irreducible totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ are either unitary or DT representations (Section 3.2), and then consider the case of reducible totally elliptic representations (Section 3.3).
+3.2. Irreducible representations
+We start with the case of irreducible surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ which are, by definition, representations whose image is not conjugate to a subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ made of upper triangular elements.
+Proposition 3.1.
+Let $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ be a reduced totally elliptic surface group representation.
+If $\rho$ is irreducible, then either $\rho$ is unitary, or it is conjugate to the composition of a DT representation with the inclusion $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}\subset\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$.
+Proof.
+For every $\gamma\in\pi_{1}\Sigma$, choose a representative $A_{\gamma}\in\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ of $\rho(\gamma)$ and write $t_{\gamma}$ for the trace of $A_{\gamma}$.
+Since $\rho$ is totally elliptic, for every $\sigma\in\pi_{1}\Sigma$ that represents a simple closed curve on $\Sigma$, the number $t_{\sigma}$ is real.
+Using trace relations for $2\times 2$ matrices, it is actually possible to write every trace $t_{\gamma}$ as a polynomial expression in the traces $t_{\sigma}$, where $\sigma$ ranges over fundamental group elements representing simple closed curves.
+This was initially observed by Goldman–Xia in their work on ergodic dynamics on character varieties [GX11, Section 2] as a consequence of classical results by Procesi [Pro76].
+In particular, this means that the traces $t_{\gamma}$ are real numbers for all $\gamma\in\pi_{1}\Sigma$.
+Since $\rho$ is irreducible, it implies that the image of $\rho$ is conjugate into a real form of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ (see for instance [MR03, Corollary 3.2.5]).
+The real forms of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ being $\operatorname{PSU}(2)$ and $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, we conclude that $\rho$ is either unitary or conjugate to a representation into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$.
+In the latter case, since $\rho$ is reduced, Theorem A applies and shows that $\rho$ is conjugate to the composition of a DT representation with the inclusion $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}\subset\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$.
+∎
+3.3. Reducible representations
+A reducible surface group representation $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ is conjugate to a representation whose image is contained in the subgroup of upper triangular elements of $\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$.
+So, up to conjugating a reducible representation $\rho$, we may decompose it into a linear part $\lambda\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}^{\times}$ and a function $z\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}$ by writing
+$\rho(\gamma)=\pm\begin{pmatrix}\lambda(\gamma)&\lambda(\gamma)^{-1}z(\gamma)\\ 0&\lambda(\gamma)^{-1}\end{pmatrix}.$
+The function $z$ turns out to be a cocycle in the sense that $z(\gamma_{1}\gamma_{2})=z(\gamma_{1})+\lambda(\gamma_{1})^{2}z(\gamma_{2})$ for every $\gamma_{1},\gamma_{2}\in\pi_{1}\Sigma$.
+If $\rho$ is totally elliptic, then $|\lambda(\gamma)|=1$ for every $\gamma\in\pi_{1}\Sigma$ because any system of geometric generators of $\pi_{1}\Sigma$ is made of simple closed curves (as the one illustrated on Figure 1).
+Proposition 3.2.
+Consider a surface $\Sigma$ of genus $g\geq 1$ and with $n\geq 0$ punctures, and let $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ be a reduced totally elliptic representation.
+If $\rho$ is reducible, then the image of $\rho$ is conjugate to a subgroup made of diagonal elements in $\operatorname{PSU}(2)$.
+Proof.
+Up to conjugating $\rho$, we may assume it is upper triangular.
+A simple computation shows that the commutator of two upper triangular $2\times 2$ matrices is elliptic if and only if it is trivial.
+More precisely, it holds that
+(3.1) $\left[\begin{pmatrix}\lambda(\gamma_{1})&\lambda(\gamma_{1})^{-1}z(\gamma_{1})\\ 0&\lambda(\gamma_{1})^{-1}\end{pmatrix},\begin{pmatrix}\lambda(\gamma_{2})&\lambda(\gamma_{2})^{-1}z(\gamma_{2})\\ 0&\lambda(\gamma_{2})^{-1}\end{pmatrix}\right]=\begin{pmatrix}1&z(\gamma_{1}\gamma_{2})-z(\gamma_{2}\gamma_{1})\\ 0&1\end{pmatrix}.$
+This statement is an analogue of Lemma 2.1.
+It means that we can apply the arguments of Section 2.2 again to conclude that $\rho$ is an abelian representation.
+Now, if $\rho$ is abelian and upper triangular, then it is conjugate to the representation $\rho^{\prime}\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ given by
+$\rho^{\prime}(\gamma)=\pm\begin{pmatrix}\lambda(\gamma)&0\\ 0&\lambda(\gamma)^{-1}\end{pmatrix}.$
+Since $\rho$ is totally elliptic, it holds that $|\lambda(\gamma)|=1$ for every $\gamma\in\pi_{1}\Sigma$ and thus $\rho^{\prime}$ is unitary. ∎
+Propositions 3.1 and 3.2 together prove Theorem C. We now proceed to the proof of Theorem D.
+The reason why Theorem C does not apply to punctured spheres is because there exist reduced totally elliptic representations $\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ that are reducible but not unitary when $\Sigma$ is a sphere with $n\geq 3$ punctures.
+Before describing such examples, we start with the following result about simple closed curves on $\Sigma$.
+Lemma 3.3.
+Let $\Sigma$ be a sphere with $n\geq 1$ punctures and $c_{1},\ldots,c_{n}$ be a system of geometric generators of $\pi_{1}\Sigma$ (like those depicted on Figure 1).
+If $\gamma\in\pi_{1}\Sigma$ represents a simple closed curve on $\Sigma$, then there exist $k\geq 1$ and $1\leq i_{1}<\cdots<i_{k}\leq n$ such that either $\gamma$ or $\gamma^{-1}$ is equal to $c_{i_{1}}\cdots c_{i_{k}}$ in the abelianization of $\pi_{1}\Sigma$.
+Proof.
+The group $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ consists of all automorphisms of $\pi_{1}\Sigma$ that map each generator $c_{i}$ to a conjugate of itself.
+In particular, two fundamental group elements that are images of each other by an automorphism in $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ coincide in the abelianization of $\pi_{1}\Sigma$.
+The image of $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ inside the group of outer automorphisms of $\pi_{1}\Sigma$ is isomorphic to the pure mapping class group of $\Sigma$, which means that $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$ preserves the subset of $\pi_{1}\Sigma$ made of all fundamental group elements representing simple closed curves on $\Sigma$.
+Moreover, if $\gamma\in\pi_{1}\Sigma$ is a fundamental group element representing a simple closed curve, then the union of the $\mathrm{Aut}^{\ast}(\pi_{1}\Sigma)$-orbits of $\gamma$ and $\gamma^{-1}$ contains exactly one product of the form $c_{i_{1}}\cdots c_{i_{k}}$ for $1\leq i_{1}<\cdots<i_{k}\leq n$ and $11\leq k\leq n-$ 1.
+This concludes the proof.
+∎
+Example 3.4.
+Using the same notation as in Lemma 3.3, we consider a representation $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ of a sphere $\Sigma$ with $n\geq 3$ punctures whose image is upper triangular.
+As above, we denote the image of $\gamma\in\pi_{1}\Sigma$ by
+$\rho(\gamma)=\pm\begin{pmatrix}\lambda(\gamma)&\lambda(\gamma)^{-1}z(\gamma)\\ 0&\lambda(\gamma)^{-1}\end{pmatrix}.$
+We claim that if the linear part $\lambda\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}^{\times}$ is chosen such that
+(3.2) $|\lambda(c_{i_{1}}\cdots c_{i_{k}})|=1\quad\text{and}\quad\lambda(c_{i_{1}}\cdots c_{i_{k}})\neq\pm 1$
+for every $1\leq i_{1}<\cdots<i_{k}\leq c_{n}$ with $11\leq k\leq n-$ 1, then $\rho$ is a reduced totally elliptic representation.
+We shall now prove this claim.
+First, observe that $\rho$ is reduced because $\lambda(c_{i})\neq\pm 1$ for every $i=1,\ldots,n$.
+Now, if $\gamma\in\pi_{1}\Sigma$ represents a simple closed curve on $\Sigma$, then $\lambda(\gamma)=\lambda(c_{i_{1}}\cdots c_{i_{k}})^{\pm 1}$ for some $1\leq i_{1}<\cdots<i_{k}\leq n$ with $11\leq k\leq n-$ 1 by Lemma 3.3.
+Since by assumption $|\lambda(c_{i_{1}}\cdots c_{i_{k}})|=1$ and $\lambda(c_{i_{1}}\cdots c_{i_{k}})\neq\pm 1$, we conclude that $\rho$ is totally elliptic.
+Let us now prove that for every $\lambda$ satisfying (3.2), we can always find an assorted cocycle $z\colon\pi_{1}\Sigma\to\operatorname{\mathbb{C}}$ such that $\rho$ is not unitary.
+Since $\rho$ is upper triangular, the image of $\rho$ has a global fixed point for the action of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ on $\mathbb{CP}^{1}$, seen here as the boundary of the 3-dimensional hyperbolic space.
+This means that $\rho$ is unitary if and only if its image fixes every point on a geodesic line in the 3-dimensional hyperbolic space, which would imply that $\rho$ is abelian.
+So, if we choose the cocycle $z$ such that there exist $i\neq j$ with $z(c_{i}c_{j})\neq z(c_{j}c_{i})$, then $\rho$ is not abelian because of (3.1), hence not unitary.
+Such cocycles $z$ always exist as soon as $n\geq 3$.
+In conclusion, by picking $\lambda$ and $z$ as we just described, we obtain a reduced totally elliptic representation $\rho\colon\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ that is reducible but not unitary.
+This finishes the proof of Theorem D.